### v1

# On the arithmetic of polynomial ideals

Abstract.
This paper investigates atomic factorizations in the monoid ${\mathcal{I}}(R)$ of nonzero ideals of a multivariate polynomial ring $R$, under ideal multiplication. Building on recent advances in factorization theory for unit-cancellative monoids, we extend techniques from [21] to construct new families of atoms in ${\mathcal{I}}(R)$, leading to a deeper understanding of its arithmetic. We further analyze the submonoid ${\mathcal{M}{\rm on}}(R)$ of monomial ideals, deriving arithmetic properties and computing sets of lengths for specific classes of ideals. The results advance the extensive study of ideal monoids within a classical algebraic framework.
Key words and phrases:
Monoids of ideals, polynomial ideal theory, monomial ideals, Krull domains, weakly Krull domains, atoms, sets of lengths, elasticity
2020 Mathematics Subject Classification:
13A15, 13B25, 13F05, 13F20, 20M12, 20M13
This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/DOC183 and 10.55776/ PAT9756623]. L. C. also acknowledges the European Union’s Horizon 2020 program (Marie Sklodowska-Curie grant 101021791) for the financial support on the early stages of the project. L. C. is a member of the National Group for Algebraic and Geometric Structures and their Applications (GNSAGA), a department of the Italian Mathematics Research Institute (INdAM)
1. Introduction
By the Fundamental Theorem of Arithmetic, every positive integer admits a unique factorization into prime numbers, up to units. This result, first formulated in Euclid’s Elements, was later found not to extend to more general settings. In particular, when considering the ring of integers of a number field (a number ring, for short), factorizations into prime elements may fail to be unique. Initially regarded as a pathological exception, starting with Narkiewicz’s work in the 1960s, this phenomenon became a subject of systematic study. The resulting field, now known as factorization theory, investigates the structure of non-unique factorizations. Its main object of study is the following notion, suitably generalizing that of factorization into primes in (multiplicative monoids of) number rings: an (atomic) factorization of an element of a commutative cancellative monoid is, up to units, a finite, unordered sequence of atoms (i.e., non-unit elements which are not the product of two non-units) whose product is equal to the element. We refer the reader to the monograph [20] for a comprehensive overview of the subject.
In recent years, while the classical theory of factorization has continued to thrive (see, for instance, [9, 11, 10, 25, 26, 27, 19]), several attempts have been made to expand its scope, considering non-commutative (see [4, 7, 35, 6]) and non-cancellative monoids (see [3, 12, 36]), even multiplicative monoids of rings with zero divisors, as in [5]. In the latter direction, the definition of unit-cancellativity was introduced in [15] as a weakening of the usual cancellativity hypothesis, and a substantial portion of the classical theory has since been developed in this more general setting; we refer [24] for a gentle introduction to the topic.
We aim to study atomic factorizations in the (commutative, unit-cancellative) monoid of nonzero ideals of the polynomial ring $K[X_{1},\dots,X_{N}]$, where $K$ is an arbitrary field, under ideal multiplication. We assume throughout the paper that $N\geq 2$, as $K[X]$ is a PID and the factorizations of its ideals correspond to the factorizations of their generating polynomials.
Our work contributes to the extensive literature on monoids of ideals: we refer to [29] and the monograph [17] for an overview, and to [33, 1, 30, 31] for more recent develpoments. A common approach in the literature is to focus on specific classes of ideals within a given ring, or semigroup ideals within a given monoid. Among these, divisorial ideals are perhaps the most prominent example. By suitably choosing a ring (or monoid) and a class of its ideals, one can obtain a fairly well-behaved monoid of ideals, and in some cases one is able to completely determine its structure. As an example, we recall two well-known results: a domain $R$ is Dedekind if and only if all its nonzero ideals are invertible, and this happens only if the monoid of all nonzero ideals is factorial, and a domain $R$ is Krull if and only if the monoid of its invertible divisorial ideals is factorial.
Compared to the “classical” examples of monoids of ideals, the structure of atomic factorizations in the monoid ${\mathcal{I}}(R)$ consisting of all nonzero ideals of a ring $R$, remains poorly understood in general. This is true even in seemingly simple cases, such as when $R$ is a polynomial ring over a field, where surprisingly little is currently known.
In the paper [21], a first step was made in the study of the arithmetic properties of ${\mathcal{I}}(R)$ when $R$ a multivariate polynomial ring. In particular, the results collected in [21, Theorem 5.1] reveal a noticeable similarity between the arithmetic of $\mathcal{I}(R)$ and that of Krull monoids with infinite class group and prime divisors in each class, despite ${\mathcal{I}}(R)$ not even being transfer Krull. More specifically, ${\mathcal{I}}(R)$ is fully elastic and each of its unions of sets of lengths is the whole set $\mathbb{N}_{\geq 2}$.
Building on these initial results, we refine the demonstrative techniques and adapt them to broader contexts. Following a review of the necessary preliminaries in Section 2, we extend the methods introduced in [21] to find several new families of atoms in ${\mathcal{I}}(R)$. In particular, we highlight in in Section 3.1 an interesting class built from sum-free subsets of $\mathbb{N}$, which includes some previously known examples.
Motivated by the problem of computing the length sets of certain ideals (see Theorem 3.7), in Section 4 we focus on the submonoid ${\mathcal{M}{\rm on}}(R)\subseteq{\mathcal{I}}(R)$ of monomial ideals of $R$ (i.e., ideals generated by monomials). We first provide some arithmetic results in the spirit of [21, Theorem 5.1], and then we determine several classes of atoms in this monoid, allowing us to compute the set of lengths of the families of ideals introduced in Theorem 3.7. The main technical ingredient which makes the study of factorizations in ${\mathcal{M}{\rm on}}(R)$ easier than in ${\mathcal{I}}(R)$ is Lemma 4.1 (see also Remark 4.4) greatly constricting the possible generators of the monomial ideals occurring in a given factorization.
We conclude the paper by outlining, in Section 5, possible directions for future research into atomic factorizations in the monoids ${\mathcal{I}}(R)$ and ${\mathcal{M}{\rm on}}(R)$, with the present work serving as a foundation.
2. Notation and preliminary results
Throughout the paper, $\mathbb{N}$ and $\mathbb{N}^{+}$ denote the sets of nonnegative and positive integers, respectively. For $a,b\in\mathbb{N}\cup\{\infty\}$ we let $\llbracket a,b\rrbracket:=\allowbreak\{x\in\mathbb{N}\colon a\leq x\leq b\}$ be the discrete interval from $a$ to $b$. For sets $A,B\subseteq\mathbb{N}$, we denote by $A+B:=\{a+b:a\in A,b\in B\}$ their sumset, by $\Delta(A)\subseteq\mathbb{N}^{+}$ the set of distances of $A$, i.e., the set of all $d\in\mathbb{N}^{+}$ such that $A\cap[a,a+d]=\{a,a+d\}$ for some $a\in A$, and by $\rho(A):={\sup A}/{\min A}\in\mathbb{Q}_{\geq 1}\cup\{\infty\}$ the elasticity of $A$ (by convention, we set $\rho(\{0\})=1$).
2.1. Factorization theory
In order to make the paper as self-contained as possible, we recall the basic notions of factorization theory. For further details, examples, and references, the reader can consult the already mentioned monograph [20].
Let $H$ be a commutative semigroup with an identity element, written multiplicatively. We denote by $H^{\times}$ the subgroup of invertible elements of $H$, also called units of $H$. If $H^{\times}=\{1\}$, we say that $H$ is reduced. An element $h\in H$ is cancellative if $h^{\prime},h^{\prime\prime}\in H$ and $hh^{\prime}=hh^{\prime\prime}$ implies $h^{\prime}=h^{\prime\prime}$, while it is unit-cancellative if $u\in H$ and $h=hu$ implies $u\in H^{\times}$. Moreover, we say that $H$ is cancellative (respectively, unit-cancellative) if all its elements are cancellative (respectively, unit-cancellative).
In this paper, a monoid is always a commutative unit-cancellative semigroup with an identity element. Note that a subsemigroup of a monoid inherits its unit-cancellativity; the subsemigroups of monoids we consider will always contain the identity element, and we will therefore speak of submonoids. A monoid $H$ is cancellative if and only if it has a quotient group, which we denote by $\mathbf{q}(H)$.
Let $H$ be a cancellative monoid. An ideal of $H$ is a subset $J\subseteq H$ such that, for any $j\in J$, $jH\subseteq J$; they are also commonly referred to as $s$-ideals (semigroup ideals), especially when $H$ also carries a ring structure, to distinguish them from usual ring ideals. For ideals $J,K$ of $H$, denote by $(J:K)$ the ideal quotient $\{g\in\mathbf{q}(H):gK\subseteq J\}$. An ideal $J$ of $H$ is divisorial if $J=(H:(H:J))$. We say that $H$ is a Mori monoid if it satisfies the ascending chain condition on divisorial ideals. Furthermore, we denote by $\hat{H}:=\{g\in\mathbf{q}(H):hg^{n}\in H\text{\ for all\ }n\in\mathbb{N}\text{\ for some\ }h\in H\}$ the complete integral closure of $H$, and we say that $H$ is completely integrally closed if $H=\hat{H}$. Finally, we say that $H$ is a Krull monoid if it is Mori and completely integrally closed.
Let $H$ be a monoid, not necessarily cancellative. A submonoid $S\subseteq H$ is divisor-closed if, whenever $h\in S,h^{\prime}\in H$ and $h^{\prime}\mid h$, we have $h^{\prime}\in S$. For a subset $A\subseteq H$, we denote by $\llbracket A\rrbracket$ the smallest divisor-closed submonoid generated by $A$. A monoid homomorphism $\varphi:H\to H^{\prime}$, where $H^{\prime}$ is cancellative, is a divisor homomorphism if $h,h^{\prime}\in H$ and $\varphi(h)\mid\varphi(h^{\prime})$ in $H^{\prime}$ implies that $h\mid h^{\prime}$ in $H$.
An element $h\in H$ is prime if $h\notin H^{\times}$ and $h\mid h^{\prime}h^{\prime\prime}$ with $h^{\prime},h^{\prime\prime}\in H$ implies $h\mid h^{\prime}$ or $h\mid h^{\prime\prime}$. An element $h\in H$ is an atom if $h\notin H^{\times}$ and $h=h^{\prime}h^{\prime\prime}$ with $h^{\prime},h^{\prime\prime}\in H$ implies $h^{\prime}\in H^{\times}$ or $h^{\prime\prime}\in H^{\times}$; the set of all atoms of $H$ is denoted by $\mathscr{A}(H)$. A factorization of an element $a\in H$ is, up to units, a finite unordered sequence of atoms of $H$, say $u_{1},\dots,u_{n}$, such that $a=u_{1}\dots u_{n}$. The length of a factorization $z$ is the number of atoms occurring in it, and it is denoted by $\lvert z\rvert$. We denote by
  • $\mathsf{Z}_{H}(a)=\mathsf{Z}(a)$ the set of factorizations of $a\in H$;
  • $\mathsf{L}_{H}(a)=\mathsf{L}(a):=\{\lvert z\rvert:z\in\mathsf{Z}(a)\}\subseteq\mathbb{N}$ the set of lengths of $a$;
  • $\mathcal{L}(H):=\{\mathsf{L}(a):a\in H\}$ the system of sets of lengths of $H$;
  • $\rho(H):=\sup\{\rho(L)\colon L\in\mathcal{L}(H)\}$ the elasticity of $H$.
Note that, if $S\subseteq H$ is a divisor-closed submonoid and $a\in S$, then $\mathsf{Z}_{S}(a)=\mathsf{Z}_{H}(a)$ and $\mathsf{L}_{S}(a)=\mathsf{L}_{H}(a)$. We say that $H$ is
  • atomic if $\mathsf{L}(a)$ is nonempty for every $a\in H$;
  • a $\mathsf{BF}$-monoid if $\mathsf{L}(a)$ is finite and nonempty for every $a\in H$;
  • an $\mathsf{FF}$-monoid if $\mathsf{Z}(a)$ is finite and nonempty for every $a\in H$;
  • half-factorial if $\lvert\mathsf{L}(a)\rvert=1$ for every $a\in H$;
  • factorial if $\lvert\mathsf{Z}(a)\rvert=1$ for every $a\in H$;
  • locally finitely generated if $\llbracket a\rrbracket$} is finitely generated up to units, for every $a\in H$;
  • fully elastic if, for any $q\in\mathbb{Q}$ with $1<q<\rho(H)$, there is an $L\in\mathcal{L}(H)$ such that $\rho(L)=q$.
Every Mori monoid is a $\mathsf{BF}$-monoid ([20, Theorem 2.2.5.1]), Krull monoids are locally finitely generated and locally finitely generated monoids are in turn $\mathsf{FF}$-monoids ([20, Prop. 2.7.8]). We note that sets of elasticities of locally finitely generated monoids have been studied in great detail in [37].
A monoid homomorphism $\theta:H\to K$ is called a transfer homomorphism if $K=\theta(H)K^{\times}$, $\theta^{-1}(K^{\times})=H^{\times}$, and, whenever $a\in H$, $\beta,\gamma\in K$ are such that $\theta(a)=\beta\gamma$, there exist $b,c\in H$ such that $a=bc$, $\theta(b)\in\beta K^{\times}$ and $\theta(c)\in\gamma K^{\times}$. In Section 4 we will use the following elementary result.
Lemma 2.1.
If $\theta:H\to K$ is a transfer homomorphism, then $\mathsf{L}_{H}(a)=\mathsf{L}_{K}(\theta(a))$ for any $a\in H$. In particular, an element $a\in H$ is an atom of $H$ if and only if $\theta(a)$ is an atom of $K$.
A monoid $H$ is transfer Krull if there exist a Krull monoid $K$ and a transfer homomorphism $\theta:H\to K$. Clearly, any Krull monoid is transfer Krull, since the identity is a transfer homormophism. Conversely, by Lemma 2.1 all transfer Krull monoids are $\mathsf{BF}$-monoids, but they need neither be Mori nor completely integrally closed, as noted in [24, Example 5.4]. All transfer Krull monoids are fully elastic, by [23, Theorem 3.1]. Some notable examples of monoids which are not transfer Krull are the integer-valued polynomials, by [18, Remark 12], and $\mathcal{P}_{\rm fin}(\mathbb{N})$, by [16, Prop. 4.12].
2.2. Multivariate polynomial rings and their ideals
Let $D$ be an integral domain with quotient field $K$, and let $R=D[X_{1},\dots,X_{N}]$ and $S=K[X_{1},\dots,X_{N}]$ be the polynomial rings in $N\geq 2$ indeterminates $X_{1},\dots,X_{N}$ over $D$ and $K$ respectively. For $m=(m_{1},\dots,m_{N})\in\mathbb{N}^{N}$, let $\mathbf{X}^{m}:=X_{1}^{m_{1}}\cdots X_{N}^{m_{N}}$ and set, for every $t\geq 0$,
$R_{t}:=\{\sum_{m\in\mathbb{N}^{N}}r_{m}\mathbf{X}^{m}\colon r_{m}\in D\text{ and }m_{1}+\cdots+m_{N}=t\},$
$S_{t}:=\{\sum_{m\in\mathbb{N}^{N}}q_{m}\mathbf{X}^{m}\colon q_{m}\in K\text{ and }m_{1}+\cdots+m_{N}=t\}.$
Then
$R=\bigoplus_{t\geq 0}R_{t}\text{ and }S=\bigoplus_{t\geq 0}S_{t}$
are graded rings, equipped with the standard grading. It is clear that $R_{t}\subseteq S_{t}$ for every $t$, that $R_{0}=D$ and $S_{0}=K$, and that $\deg X_{1}=\dots=\deg X_{N}=1$. It follows that every $f\in S$ can be uniquely written as $f=\sum_{t\geq 0}f_{t}$, with $f_{t}\in S_{t}$ for every $t$.
We can then define the min-degree of a nonzero polynomial $f\in S$ as the smallest nonnegative integer $d$ such that $f_{d}$ is nonzero, and we denote it as $\operatorname{\mathsf{m-deg}}(f)$. We set $\operatorname{\mathsf{m-deg}}(0):=+\infty$. For all $f,g\in S\setminus\{0\}$, the following properties hold:
  1. (a)
    $\operatorname{\mathsf{m-deg}}(fg)=\operatorname{\mathsf{m-deg}}(f)+\operatorname{\mathsf{m-deg}}(g)$;
  2. (b)
    $\operatorname{\mathsf{m-deg}}(f+g)\geq\min\{\operatorname{\mathsf{m-deg}}(f),\operatorname{\mathsf{m-deg}}(g)\}$, with equality if $\operatorname{\mathsf{m-deg}}(f)\neq\operatorname{\mathsf{m-deg}}(g)$.
Let $I\subseteq R$ be an ideal of $R$. We define the min-degree $\operatorname{\mathsf{m-deg}}(I)$ of $I$ as the smallest nonnegative integer $d$ such that $I$ contains a polynomial whose min-degree is $d$. The min-degree of the zero ideal is set equal to $\infty$.
Remark 2.2.
It is easy to see from the properties of the min-degree, that if $I$ is a nonzero finitely generated ideal of $R$, i.e., $I=\langle g_{1},\dots,g_{n}\rangle$ with $n\geq 1$ and $g_{1},\dots,g_{n}\in R$, then
$\operatorname{\mathsf{m-deg}}(I)=\min\{\operatorname{\mathsf{m-deg}}(g_{i})\colon i\in\llbracket 1,n\rrbracket\}.$
Let ${\mathcal{I}}(R)$ be the semigroup of nonzero (ring) ideals of $R$, with the usual ideal multiplication. This semigroup is well known (see, e.g., [21, Prop. 2.1 (4)]) to be reduced and unit-cancellative, and thus a monoid in the sense specified in Section 2.1, being obviously commutative.
Remark 2.3.
It is of interest to understand how far is ${\mathcal{I}}(R)$ from being cancellative. We recall here a remarkable characterization of the cancellative elements of ${\mathcal{I}}(R)$ due to Anderson and Roitman, valid for any commutative ring $R$, not necessarily even a domain: an ideal $I$ of $R$ is cancellative if and only if it is a regular principal ideal at each localization at a maximal ideal [2]. If we consider $R=D[X_{1},\dots,X_{N}]$, with $D$ a noetherian domain and $N\geq 2$, it is easy to check that the cancellative elements of ${\mathcal{I}}(R)$ are precisely the principal ideals of $R$.
It was proved in Lemma 5.5 of [21] that the map $\operatorname{\mathsf{m-deg}}\colon\mathcal{I}(R)\to\mathbb{N}\text{ defined by }I\mapsto\operatorname{\mathsf{m-deg}}(I)$ is a semigroup homomorphism, i.e., for every $I,J\in\mathcal{I}(R)$,
$\operatorname{\mathsf{m-deg}}(IJ)=\operatorname{\mathsf{m-deg}}(I)+\operatorname{\mathsf{m-deg}}(J).$ (1)
For a nonzero ideal $I$ of $R$ and a nonnegative integer $i$, we let
$I[i]:=\{f_{i}\colon f\in I\}$
be the set of the homogeneous components of degree $i$ of the elements of $I$, which is a $D$-module contained in $R_{i}$. For $I,J\in{\mathcal{I}}(R)$ and $i,j\in\mathbb{N}$,
$I[i]\cdot J[j]:=\left\{\sum_{s=1}^{k}x_{s}y_{s}\colon k\geq 1,x_{s}\in I[i],y_{s}\in J[j]\right\}$
is also a $D$-module. Moreover, we let
$I_{K}[i]:={\rm Span}_{K}\{I[i]\}$
be the $K$-linear span of the elements of $I[i]$. It is clear that $I_{K}[i]$ is a finite dimensional $K$-vector space contained in $S_{i}$.
Remark 2.4.
If $I=\langle g_{1},\dots,g_{n}\rangle$ is a nonzero finitely generated ideal of $R$ and $\operatorname{\mathsf{m-deg}}(I)=m\in\mathbb{N}$, then
$I_{K}[m]={\rm Span}_{K}\{(g_{i})_{m}\colon\operatorname{\mathsf{m-deg}}(g_{i})=m\}.$
Lemma 2.5 (Lemma 5.6 of [21]).
Let $I,J\in\mathcal{I}(R)$ and set $\operatorname{\mathsf{m-deg}}(I)=d$, $\operatorname{\mathsf{m-deg}}(J)=e$ and $\operatorname{\mathsf{m-deg}}(IJ)=m$. The following hold:
(i) $(IJ)[m]=I[d]\cdot J[e]$;
(ii) $(IJ)_{K}[m]=I_{K}[d]\cdot J_{K}[e]$.
Remark 2.6.
Throughout the rest of the paper, we will consider ideals of $R$ generated by polynomials in two indeterminates, say $X_{1}$ and $X_{2}$. To slightly ease the notation, we will rename these indeterminates to $X$ and $Y$ respectively. Note that if an ideal is generated only by polynomials in $X$ and $Y$, ideals occurring in any of its factorizations are necessarily also generated by polynomials in $X$ and $Y$ only. We will use this fact in the rest of the paper without further explanation.
Finally, let $\leq$ be the lexicographic order on monomials of $S$ defined by $X>Y>X_{3}>\dots>X_{N}$; we denote by $\operatorname{\mathsf{in}}(f)$ the initial monomial of $f\in S$ with respect to $\leq$, that is, the largest monomial occurring in $f$ with a non-zero coefficient.
2.3. A link between polynomial ideals and power monoids
Keeping the notation of the previous subsection, we let $R=D[X_{1},\dots,X_{N}]$, with $D$ an integral domain, be a polynomial ring in $N\geq 2$ indeterminates. As mentioned in the introduction, the study of ${\mathcal{I}}(R)$ from the perspective of factorization theory was initiated [21, Sect. 5], where it was shown that when ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid, it is fully elastic and each of its unions of sets of lengths equals $\mathbb{N}_{\geq 2}$ [21, Theorem 5.1].
The proof of the main theorem relies on the construction of several families of atoms in $\mathcal{I}(R)$:
$\mathfrak{b}_{i}(X,Y):=\langle X^{i},Y^{i}\rangle\ {\rm for}\ i\in\mathbb{N}^{+},$
$\mathfrak{c}_{2i+1}(X,Y):=\langle\{X^{2i+1},X^{2i}Y\}\cup\{X^{2i-j}Y^{j+1}\colon j\in\llbracket 1,2i+1\rrbracket\cap 2\mathbb{N}\}\rangle\ {\rm for}\ i\in\mathbb{N}^{+},$
$\mathfrak{c}_{2i}(X,Y):=\langle\{X^{2i},X^{2i-1}Y\}\cup\{X^{2i-j}Y^{j}\colon j\in\llbracket 1,2i\rrbracket\cap 2\mathbb{N}\}\rangle\ {\rm for}\ i\in\mathbb{N}_{\geq 3}.$
These families of atoms share a common structure: as explained in [21, Example 5.13], they correspond to families of atoms in the monoid of finite nonempty subsets of $\mathbb{N}$ endowed with set-wise addition, known as the (finitary) power monoid of $\mathbb{N}$ and denoted by $\mathcal{P}_{{\rm fin}}(\mathbb{N})$. More precisely, $\mathcal{I}(R)$ has a submonoid which is isomorphic to $\mathcal{P}_{{\rm fin}}(\mathbb{N})$. Indeed, the map
$\Phi\colon\ \mathcal{P}_{{\rm fin}}(\mathbb{N})\to{\mathcal{I}}(R),\ A\mapsto I_{A}$ (2)
that sends a set $A\in\mathcal{P}_{{\rm fin}}(\mathbb{N})$, say $A=\{n_{1},\dots,n_{\ell}\}$ with $\ell\geq 1$ and $0\leq n_{1}<\dots<n_{\ell}$, to the ideal
$I_{A}:=\langle X^{n_{\ell}-n_{1}}Y^{n_{1}},X^{n_{\ell}-n_{2}}Y^{n_{2}},\dots,X^{n_{\ell}-n_{\ell-1}}Y^{n_{\ell-1}},Y^{n_{\ell}}\rangle,$
is readily seen to be an injective monoid homomorphism. The image of this map, denoted by $\mathcal{M}_{2}(X,Y)$, consists precisely of those ideals which can be generated by $Y^{m}$ and a (possibly empty) set of monomials in $X$ and $Y$ of degree $m$, for some $m\in\mathbb{N}$. In particular, $\mathcal{M}_{2}(X,Y)\subseteq{\mathcal{M}{\rm on}}(R)\subseteq{\mathcal{I}}(R)$, where ${\mathcal{M}{\rm on}}(R)$ is the monoid of monomial ideals of $R$ (i.e., ideals generated by monomials).
The study of the arithmetic properties of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$ was started by Fan and Tringali in [16]. In particular, it follows from [16, Prop.s 3.2 and 3.5] that $\mathcal{P}_{{\rm fin}}(\mathbb{N})$ is a commutative, reduced, unit-cancellative semigroup with $\{0\}$ as the identity element. Among their results, Fan and Tringali show that for any $i\in\mathbb{N}^{+}$, the sets $\{0,i\}$ and $\{0,1,3,\dots,2i+1\}$ are atoms, while $\{0,1,2,\dots,2i\}$ is an atom for $i\in\mathbb{N}_{\geq 2}$ (see [16, Prop.s 4.1(iv) and 4.2]). The correspondence $\Phi$ in (2) sends these sets to the ideals $\mathfrak{b}_{i}(X,Y)$, $\mathfrak{c}_{2i+1}(X,Y)$, and $\mathfrak{c}_{2i}(X,Y)$, respectively. In [21, Prop. 5.10], it is proved that $\mathfrak{b}_{i}(X,Y)$ and $\mathfrak{c}_{2i+1}(X,Y)$ are atoms of $\mathcal{I}(R)$ for any $i\in\mathbb{N}^{+}$, while $\mathfrak{c}_{2i}(X,Y)$ is an atom only for $i\geq 3$. Note that, a priori, the above correspondence ensures only that these ideals are atoms of the submonoid $\mathcal{M}_{2}(X,Y)$, as monoid isomorphisms are clearly atom-preserving. Indeed, it turns out that in general $\Phi$ does not send atoms of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$ to atoms of $\mathcal{I}(R)$, as the set $C=\{0,1,2,4\}$ is easily seen to be an atom of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$, while, when $2$ is a unit of $D$, (see [21, Remark 5.11])
$I_{C}=\mathfrak{c}_{4}(X,Y)=\langle X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}\rangle=\langle X^{2},XY+Y^{2}\rangle\langle X^{2},XY-Y^{2}\rangle.$
On the other hand, we will show in Remark 4.3 that $I_{C}$ is an atom of ${\mathcal{M}{\rm on}}(R)$. Thus, it might be the case that the map $\Phi$ sends atoms of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$ to atoms of ${\mathcal{M}{\rm on}}(R)$.
3. Atoms and sets of lengths in ${\mathcal{I}}(R)$
Throughout this section, we maintain the notations introduced in Section 2.2, and assume that ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid. This assumption is clearly satisfied if, for instance, $D$ is noetherian (in particular, if $D=K$). Our aim is to find new classes of atoms in $\mathcal{I}(R)$, by extending the methods introduced in [21, Sect. 5] to further families of ideals of $R$, corresponding (via the map $\Phi$ in (2) from $\mathcal{P}_{\rm fin}(\mathbb{N})$ to ${\mathcal{I}}(R)$) to atoms of the reduced power monoid $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ of $\mathbb{N}$, i.e., the monoid of nonempty finite subsets of $\mathbb{N}$ containing 0 with setwise addition. Note that $\mathscr{A}(\mathcal{P}_{\rm fin}(\mathbb{N}))=\mathscr{A}(\mathcal{P}_{{\rm fin},0}(\mathbb{N}))\cup\{\{1\}\}$, since every finite set of positive integers $A=\{a_{1},\dots,a_{n}\}$ with $n\geq 2$ or $a_{1}\geq 2$ can be written as $A=(A-1)+\{1\}$. It is also worth noting that the monoid homomorphism $\Phi$ restricts to a monoid isomorphism between $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ and the ideals of $R$ generated, for some $m\in\mathbb{N}$, by $X^{m}$, $Y^{m}$ and a set of monomials in $X$ and $Y$ of degree $m$ (cf. [21, Example 5.13]).
The following lemma, generalizing the arguments used in [21], collects a number of necessary conditions for the existence of a nontrivial factorization for an ideal $I_{A}$, where $A\in\mathcal{P}_{{\rm fin},0}(\mathbb{N})$. They will be a key tool in the proofs of the main results of this section, namely, Theorem 3.3 and Theorem 3.7.
Lemma 3.1.
Let $m\in\mathbb{N}^{+}$, $\mathcal{N}$ be a set of monomials of degree $m$ in $X$ and $Y$, and $I=\langle X^{m},Y^{m},\mathcal{N}\rangle$ be an ideal of $R$. Suppose that $I=\mathfrak{a}\cdot\mathfrak{b}$, for some $\mathfrak{a},\mathfrak{b}\in\mathcal{I}(R)$ such that $I\subsetneq\mathfrak{a}$, $\mathfrak{b}\subsetneq R$, $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$, $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$, ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$. Then, the following hold.
(i) m=d+em=d+eand $I_{K}[m]=\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e]$.
(ii) $d\geq 1,e\geq 1$.
(iii) $\mathfrak{a}_{K}[d]$has dimension $r\geq 2$ and admits a $K$-basis $\{f_{1},\dots,f_{r}\}$ with $X^{d}=\operatorname{\mathsf{in}}(f_{1})>\operatorname{\mathsf{in}}(f_{2})>\dots>\operatorname{\mathsf{in}}(f_{r})$. Analogously, $\mathfrak{b}_{K}[e]$ has dimension $r\geq 2$ and admits a $K$-basis $\{g_{1},\dots,g_{s}\}$ with $X^{e}=\operatorname{\mathsf{in}}(g_{1})>\operatorname{\mathsf{in}}(g_{2})>\dots>\operatorname{\mathsf{in}}(g_{s})$.
(iv) For every $j\in\llbracket 2,r\rrbracket$ there exists $a_{j}\in\llbracket 1,d\rrbracket$ such that $\operatorname{\mathsf{in}}(f_{j})=X^{d-a_{j}}Y^{a_{j}}$. Analogously, for each $k\in\llbracket 2,s\rrbracket$ there exists $b_{k}\in\llbracket 1,e\rrbracket$ such that $\operatorname{\mathsf{in}}(g_{k})=X^{e-b_{k}}Y^{b_{k}}$.
(v) Set $a_{1}=b_{1}=0$; then, for every $j\in\llbracket 1,r\rrbracket$ and every $k\in\llbracket 1,s\rrbracket$ we have $X^{m-(a_{j}+b_{k})}Y^{a_{j}+b_{k}}\in\{X^{m},Y^{m}\}\cup\mathcal{N}.$
Proof.
It follows from Remark 2.2 that $\operatorname{\mathsf{m-deg}}(I)=m$, and item (i) then follows from equation (1) and Lemma 2.5(i); item (ii) follows from [21, Lemma 5.8], while item (iv) is a consequence of item (iii). We are thus left to prove items (iii) and (v).
Since ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$, we get from [21, Lemma 5.9] that there are bases $\{f_{1},\dots,f_{r}\}$ of $\mathfrak{a}_{K}[d]$ and $\{g_{1},\dots,g_{s}\}$ of $\mathfrak{b}_{K}[e]$ with $\operatorname{\mathsf{in}}(f_{1})>\operatorname{\mathsf{in}}(f_{2})>\dots>\operatorname{\mathsf{in}}(f_{r})$ and $\operatorname{\mathsf{in}}(g_{1})>\operatorname{\mathsf{in}}(g_{2})>\dots>\operatorname{\mathsf{in}}(g_{s})$. Note that the $\operatorname{\mathsf{in}}(f_{j})$ and $\operatorname{\mathsf{in}}(g_{k})$ are monomials of degree $d$ and $e$ respectively, for all $j\in\llbracket 1,r\rrbracket$ and $k\in\llbracket 1,s\rrbracket$. If $r=1$, we get from (i) that
$I_{K}[m]\ni X^{m}=\sum_{k=1}^{s}\alpha_{k}f_{1}g_{k}\ \quad{\rm for\ some}\ \alpha_{1},\dots.\alpha_{s}\in K.$
Thus, we obtain $f_{1}=X^{d}$, implying in turn that $Y^{m}\notin{\rm Span}_{K}\{f_{1}\}\cdot\mathfrak{b}_{K}[e]$, contradicting (i). Therefore, it must be $r\geq 2$, and similarly we obtain that also $s\geq 2$. On the other hand, if $\operatorname{\mathsf{in}}(f_{1})\neq X^{d}$, then for every $f\in\mathfrak{a}_{K}[d]\setminus\{0\}$ we have $\operatorname{\mathsf{in}}(f)<X^{d}$, hence $\operatorname{\mathsf{in}}(fg)<X^{m}$ for every $f\in\mathfrak{a}_{K}[d]\setminus\{0\}$, $g\in\mathfrak{b}_{K}[e]\setminus\{0\}$, and so $X^{m}\notin I_{K}[m]$, contradicting (i). Thus, $\operatorname{\mathsf{in}}(f_{1})=X^{d}$, and analogously we can prove that also $\operatorname{\mathsf{in}}(g_{1})=X^{e}$. This concludes the proof of item (iii).
To prove item (v), consider, for any $j\in\llbracket 1,r\rrbracket$ and $k\in\llbracket 1,s\rrbracket$,
$\operatorname{\mathsf{in}}(f_{j}g_{k})=\operatorname{\mathsf{in}}(f_{j})\operatorname{\mathsf{in}}(g_{k})=X^{d-a_{j}+e-b_{k}}Y^{a_{j}+b_{k}}=X^{m-(a_{j}+b_{k})}Y^{a_{j}+b_{k}},$
where we have used (iv) and (i). Since $f_{j}g_{k}\in I_{K}[m]={\rm Span}_{K}(\{X^{m},Y^{m}\}\cup\mathcal{N})$, we must also have $\operatorname{\mathsf{in}}(f_{j}g_{k})\in I_{K}[m]$, and the conclusion follows from the above equalities. ∎
3.1. Atoms constructed from sum-free sets of positive integers
We are now in the position to extend the results on the atomicity of $\mathfrak{b}_{i}(X,Y)$ and $\mathfrak{c}_{2i+1}(X,Y)$ to a larger family of ideals. Before proceeding, let us recall a classical definition from additive combinatorics: a nonempty subset $A\subseteq\mathbb{N}$ is sum-free if $A\cap(A+A)=\emptyset$. Easy examples of sum-free sets include any set of odd positive integers and any subset of $\{\lceil{n/2}\rceil,\dots,n\}$, where $n\in\mathbb{N}$ and $\lceil\ \rceil$ denotes the usual ceiling function. In fact, Green proved in [28] that these subsets account for “most” of the sum-free sets; more precisely, he proved that the number of sum-free subsets of $\{1,\dots,n\}$ is $O(2^{\nicefrac{{n}}{{2}}})$ for $n\in\mathbb{N}$, as predicted by the Cameron-Erdös Conjecture.
Clearly, a finite sum-free set is contained in $\mathcal{P}_{{\rm fin}}(\mathbb{N})\setminus\mathcal{P}_{{\rm fin},0}(\mathbb{N})$. It follows that no sum-free set different from $\{1\}$ is an atom of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$ (see the remarks at the beginning of the section). In preparation for Theorem 3.3, we note that adjoining the zero element to sum-free subsets of $\mathbb{N}$ turns them into atoms of $\mathcal{P}_{{\rm fin}}(\mathbb{N})$.
Lemma 3.2.
If $A$ is finite and sum-free, then $\{0\}\cup A$ is an atom of $\mathcal{P}_{\rm fin}(\mathbb{N})$.
Proof.
Suppose that there exist $B,C\in\mathcal{P}_{\rm fin}(\mathbb{N})\setminus\{\{0\}\}$ such that
$\{0\}\cup A=B+C.$
Firstly, note that $0\in B\cap C$, for otherwise 0 cannot be an element of $B+C$. Furthermore, since $B$ and $C$ both differ from $\{0\}$, both $\beta:=\max{B}$ and $\gamma:=\max{C}$ are positive integers. The set $B+C$ contains, in particular, the elements $\beta,\gamma$ and $\beta+\gamma$; since they are all positive, they necessarily belong to $A$, but this is a contradiction, as $A$ was supposed to be sum-free. ∎
For every $i\in\mathbb{N}^{+}$, the sets $\beta_{i}=\{i\}$ and $\delta_{2i+1}=\{2\llbracket 0,i\rrbracket+1\}$ are clearly sum-free. We then get from Lemma 3.2 an alternative proof of the fact (mentioned in the introduction and proved in [16]) that $\{0\}\cup\beta_{i}$ and $\{0\}\cup\delta_{2i+1}$ are atoms of $\mathcal{P}_{\rm fin}(\mathbb{N})$. The next theorem (the first of the main results of this section) provides a general method for constructing atoms in ${\mathcal{I}}(R)$ from sum-free subsets of $\mathbb{N}$ and yields, as a special case, the atomicity of $\mathfrak{b}_{i}(X,Y)=I_{\{0\}\cup\beta_{i}}$ and $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i+1}}$.
Theorem 3.3.
If $A$ is a finite sum-free subset of $\mathbb{N}$, then $I_{\{0\}\cup A}$ is an atom of $\mathcal{I}(R)$.
Proof.
Suppose that $I_{\{0\}\cup A}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in\mathcal{I}(R)\setminus\{R\}$. Set $m=\max{A}=\operatorname{\mathsf{m-deg}}(I_{\{0\}\cup A})$, $d=\operatorname{\mathsf{m-deg}}(\mathfrak{a})$ and $e=\operatorname{\mathsf{m-deg}}(\mathfrak{b})$. Since $\{0\}\cup A\in\mathcal{P}_{{\rm fin},0}(\mathbb{N})$, we can apply Lemma 3.1. Therefore, we obtain $d\geq 1$, $e\geq 1$, $m=d+e$,
${(I_{\{0\}\cup A})}_{K}[m]=\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e],$
$\mathfrak{a}_{K}[d]={\rm Span}_{K}\{f_{1},\dots,f_{r}\}$, where $r\geq 2$ and $X^{d}=\operatorname{\mathsf{in}}(f_{1})>\dots>\operatorname{\mathsf{in}}(f_{r})$, $\mathfrak{b}_{K}[e]={\rm Span}_{K}\{g_{1},\dots,g_{s}\}$, where $s\geq 2$ and $X^{e}=\operatorname{\mathsf{in}}(g_{1})>\dots>\operatorname{\mathsf{in}}(g_{s})$. Let $\operatorname{\mathsf{in}}(f_{2})=X^{d-a}Y^{a}$ for some $a\in\llbracket 1,d\rrbracket$ and $\operatorname{\mathsf{in}}(g_{2})=X^{e-b}Y^{b}$ for some $b\in\llbracket 1,e\rrbracket$. Then Lemma 3.1((v)) yields
$X^{m-a}Y^{a},X^{m-b}Y^{b},X^{m-(a+b)}Y^{a+b}\in\{X^{m-c}Y^{c}\colon c\in\{0\}\cup A\},$
i.e., $a,b,a+b\in\{0\}\cup A$. Since both $a$ and $b$ are positive, we actually get $a,b,a+b\in A$, but this contradicts the assumption that $A$ is a sum-free set. ∎
In light of the remarks at the beginning of this section, the above theorem gives a substantial number of examples of atoms of ${\mathcal{I}}(R)$. As a further corollary, we also obtain an $\mathcal{I}(R)$-analogue of [16, Prop. 4.1(v)].
Corollary 3.4.
Let $m\in\mathbb{N}_{\geq 2}$. For any integer $j\in\llbracket 1,m-1\rrbracket$, the ideal $I=I_{\{0,j,m\}}=\langle X^{m},X^{m-j}Y^{j},Y^{m}\rangle$ is an atom of $\mathcal{I}(R)$ if and only if $m\neq 2j$.
Proof.
If $m=2j$, then $I={\langle X^{j},Y^{j}\rangle}^{2}$ is clearly not an atom. Conversely, if $m\neq 2j$, then the set $\{j,m\}$ is sum-free, and $I=I_{\{0,j,m\}}$ is an atom by Theorem 3.3. ∎
3.2. A further family of atoms
For every $i\in\mathbb{N}_{\geq 2}$, the set $\delta_{2i}=\{1,2\llbracket 1,i\rrbracket\}$ is clearly not sum-free. Nonetheless, as recalled in Section 2.3, $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i}}$ is an atom for every $i\geq 3$. In this section, we will provide a further class of atoms not accounted for by Theorem 3.3.
Following [16, Prop. 4.10], fix a positive integer $n\geq 2$ and let $a_{1},\dots,a_{n},a_{n+1}$ be positive integers satisfying the following conditions:
  1. (C1)
    $a_{n+1}=a_{1}+\dots+a_{n-1}+2a_{n}$,
  2. (C2)
    $a_{i+1}>2(a_{1}+\dots+a_{i})$ for all $i\in\llbracket 1,n-1\rrbracket$.
Further, define
$A_{n}:=\left\{a_{I}\colon I\subseteq\llbracket 1,n-1\rrbracket\right\},$
$B_{n}:=A_{n}\cup\{a_{\llbracket 1,n\rrbracket}\}\cup(A_{n}+a_{n+1}),$
$C_{n}:=\left\{a_{I}\colon I\subseteq\llbracket 1,n+1\rrbracket\right\},$
where we set $a_{\emptyset}=0$ and $a_{I}=\sum_{i\in I}a_{i}$ for every nonempty subset $I\subseteq\llbracket 1,n+1\rrbracket$. In [16, Prop. 4.10(i)], it is proved that, for every $n\geq 2$, the set $B_{n}$ is an atom of $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$.
Remark 3.5.
By the definition of the $a_{i}$’s it is clear that $\max(B_{n})=a_{\llbracket 1,n-1\rrbracket}+a_{n+1}=a_{\llbracket 1,n-1\rrbracket\cup\{a_{n+1}\}}$. By condition (C1), if $a_{\llbracket 1,n-1\rrbracket}$ is odd (resp. even), then $a_{n+1}$ is odd (resp. even), hence $m=\max(B_{n})$ is always even. In particular, $m=a_{\llbracket 1,n-1\rrbracket}+a_{n+1}=2a_{\llbracket 1,n-1\rrbracket}+2a_{n}=2a_{\llbracket 1,n\rrbracket}$.
Let us focus our attention to the ideals of $R$
$I_{B_{n}}=\langle X^{m-b}Y^{b}\colon b\in B_{n}\rangle\text{ and }I_{C_{n}}=\langle X^{a_{\llbracket 1,n+1\rrbracket}-c}Y^{c}\colon c\in C_{n}\rangle.$
Remark 3.6.
The following identities can be easily verified by a direct computation or using [16, Prop. 4.10 (ii)] and the monoid homomorphism $\Phi\colon\ {P}_{{\rm fin}}(\mathbb{N})\to{\mathcal{I}}(R)$.
$I_{C_{n}}=\mathfrak{b}_{a_{1}}\cdot\dots\cdot\mathfrak{b}_{a_{n+1}}=\mathfrak{b}_{a_{n}}\cdot I_{B_{n}}.$ (3)
We recall here that $\mathfrak{b}_{i}=\langle X^{i},Y^{i}\rangle\in\mathscr{A}({\mathcal{I}}(R))$ for every $i\in\mathbb{N}^{+}$.
The next theorem provides a new class of atoms of ${\mathcal{I}}(R)$, namely the ideals $I_{B_{n}}$, and gives bounds on the set of lengths of the ideals $I_{C_{n}}$.
Theorem 3.7.
(1) $I_{B_{n}}$is an atom of $\mathcal{I}(R)$ for every $n\geq 2$.
(2) $\{2,n+1\}\subseteq\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})\subseteq\llbracket 2,n+1\rrbracket$.
Before proving this result, we need some technical lemmas based on the properties of the positive integers $a_{1},\dots,a_{n+1}$
Lemma 3.8.
Let $n,k$ be positive integers such that $n\geq 2$ and $k\leq n$ and let $I,J\subseteq\llbracket 1,k\rrbracket$. Then $a_{I}+a_{J}=a_{H}$ for some $H\subseteq\llbracket 1,k\rrbracket$ if and only if $I\cap J=\emptyset$.
Proof.
One of the implications is trivial: if $I\cap J=\emptyset$, then $a_{I}+a_{J}=a_{I\cup J}$. Conversely, let $I\cap J\neq\emptyset$ and assume to the contrary that there exists $H\subseteq\llbracket 1,k\rrbracket$ such that
$a_{H}=a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}=a_{(I\cup J)\setminus(I\cap J)}+2a_{I\cap J}.$ (4)
Since $a_{I\cap J}>0$, in particular $H\neq\emptyset$. Let $h:=\max H$ and $\ell:=\max(I\cup J)$. Then equality (4) and condition (C2) imply $h=\ell$ and $h\in(I\cup J)\setminus(I\cap J)$. In fact, if $h>\ell$, then (using the fact that both $h$ and $\ell$ are at most $k$) we obtain $I,J\subseteq\llbracket 1,h-1\rrbracket$ and
$a_{H}\geq a_{\{h\}}>2a_{\llbracket 1,h-1\rrbracket}\geq 2a_{(I\cup J)\setminus(I\cap J)}+2a_{I\cap J}\geq a_{I}+a_{J};$
if $\ell>h$, then $H\subseteq\llbracket 1,\ell-1\rrbracket$ and
$a_{H}<2a_{\llbracket 1,\ell-1\rrbracket}<a_{\{\ell\}}\leq a_{I\cup J}<a_{I}+a_{J};$
if $h=\ell$ and $h\in I\cap J$, then
$a_{H}\leq a_{\{h\}}+a_{\llbracket 1,h-1\rrbracket}<2a_{\{h\}}\leq 2a_{I\cap J}\leq a_{I}+a_{J}.$
Now, since $h=\ell\in(I\cup J)\setminus(I\cap J)$, it follows that $((I\cup J)\setminus(I\cap J))\setminus\{h\}=((I\cup J)\setminus\{h\})\setminus(I\cap J)$; moreover, equality (4) yields
$a_{H\setminus\{h\}}=a_{(I\cup J)\setminus\{h\}}+a_{I\cap J}=a_{((I\cup J)\setminus\{h\})\setminus(I\cap J)}+2a_{I\cap J}.$
If $H\setminus\{h\}=\emptyset$, we immediately get a contradiction. Hence, set $h^{\prime}:=\max(H\setminus\{h\})$ and $\ell^{\prime}:=\max((I\cup J)\setminus\{h\})$ (note that $(I\cup J)\setminus\{h\}\neq\emptyset$ since $I\cap J\neq\emptyset$). Arguing as before, we obtain $h^{\prime}=\ell^{\prime}\in((I\cup J)\setminus\{h\})\setminus(I\cap J)$. Iterating this argument, we arrive at the following equation
$a_{x}=a_{(I\cup J)\setminus(H\setminus\{x\})}+a_{I\cap J},$ (5)
where $x\in H$ and $x\in(I\cup J)\setminus(I\cap J)$. By subtracting $a_{x}$ from both sides of equation (5), we finally obtain $a_{I\cap J}=0$, a contradiction. ∎
Lemma 3.9.
For a positive integer $n\geq 2$ and $I,J,H\subseteq\llbracket 1,n-1\rrbracket$, assume that
$a_{\llbracket 1,n-1\rrbracket}+a_{H}=a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}.$
Then, $I\cup J=\llbracket 1,n-1\rrbracket$ and $H=I\cap J$.
Proof.
We first observe that $n-1\in I\cup J$. If $n=2$ this is trivial, otherwise, if $n>2$ and $I\cup J\subseteq\llbracket 1,n-2\rrbracket$, we get the contradiction
$a_{\llbracket 1,n-1\rrbracket}+a_{H}\geq a_{n-1}>2a_{\llbracket 1,n-2\rrbracket}\geq a_{I}+a_{J}.$
Moreover $n-1\in H$ if and only if $n-1\in I\cap J$. In fact if $n-1\in H$ and $n-1\notin I\cap J$, then
$a_{\llbracket 1,n-1\rrbracket}+a_{H}\geq 2a_{n-1}>a_{n-1}+2a_{\llbracket 1,n-2\rrbracket}\geq a_{I}+a_{J},$
while if $n-1\in I\cap J$ and $n-1\notin H$, then
$a_{\llbracket 1,n-1\rrbracket}+a_{H}\leq a_{n-1}+2a_{\llbracket 1,n-2\rrbracket}<2a_{n-1}\leq a_{I}+a_{J},$
which is a contradiction in both cases. Thus, $a_{\llbracket 1,n-1\rrbracket}+a_{H}=a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}$ implies that $a_{\llbracket 1,n-2\rrbracket}+a_{H\setminus\{n-1\}}=a_{I\setminus\{n-1\}}+a_{J\setminus\{n-1\}}=a_{(I\cup J)\setminus\{n-1\}}+a_{(I\cap J)\setminus\{n-1\}}$. Arguing recursively as in the previous lemma, we prove the statement after n1n-1 steps. ∎
Lemma 3.10.
Let $n\geq 2$. The following statements hold:
(iii) For every $I,J\subseteq\llbracket 1,n-1\rrbracket$, $a_{I}+a_{J}\in A_{n}$ if and only if $I\cap J=\emptyset$.
(iv) For every $I\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$ and $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\}$ (note that $\cup$ is a disjoint union), $a_{I}+a_{J}\in B_{n}$ if and only if $I\cap J=\emptyset$. Moreover, $a_{\llbracket 1,n\rrbracket}+a_{I}\in B_{n}$ if and only if $I=\emptyset$.
(v) For every $I,J\subseteq\llbracket 1,n+1\rrbracket$, $a_{I}+a_{J}\in C_{n}$ if and only if either $I\cap J=\emptyset$ or $I\cup J=\llbracket 1,n\rrbracket$ and $n\in I\cap J$.
Proof.
(iii) The statement follows from Lemma 3.8 with k=n1k=n-1. In fact, $a_{I}+a_{J}\in A_{n}$ if and only if there is $H\subseteq\llbracket 1,n-1\rrbracket$ such that $a_{H}=a_{I}+a_{J}$.
(iv) We first assume $I\cap J=\emptyset$. If $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$, then $a_{I}+a_{J}\in A_{n}\subseteq B_{n}$ by (iii). If $J=X\cup\{n+1\}$ for some $X\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$, then $I\cap X=\emptyset$ and, by (iii), $a_{I}+a_{X}\in A_{n}$. Thus $a_{I}+a_{J}=a_{I}+a_{X}+a_{n+1}\in A_{n}+a_{n+1}\subseteq B_{n}$.
On the other hand, let $I\subseteq\llbracket 1,n-1\rrbracket$, $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\}$ and assume $a_{I}+a_{J}\in B_{n}$. Then $a_{I}+a_{J}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, or $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$, with $X\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$. Let us first observe that $H\neq\llbracket 1,n\rrbracket$. Indeed, if we assume that $H=\llbracket 1,n\rrbracket$, if $n+1\in J$, then $a_{I}+a_{J}\geq a_{n+1}>a_{H}$, while if $n+1\notin J$ (so $I,J\subseteq\llbracket 1,n-1\rrbracket$), $a_{I}+a_{J}\leq 2a_{\llbracket 1,n-1\rrbracket}<a_{n}\leq a_{H}$.
If $H\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$, then $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$ and $a_{H}\in A_{n}$, hence $I\cap J=\emptyset$, by (iii). If $H=X\cup\{n+1\}$ for some $X\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$, then $J=Y\cup\{n+1\}$ for some $Y\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$ and $a_{I}+a_{Y}=a_{X}$. Note that $a_{X}\in A_{n}$, so, by (iii), $I\cap Y=\emptyset$ and hence $I\cap J=\emptyset$.
Finally, we prove that $a_{\llbracket 1,n\rrbracket}+a_{I}\in B_{n}$ only if $I=\emptyset$ (the reverse implication is trivial). Assume $a_{\llbracket 1,n\rrbracket}+a_{I}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$, with $X\subseteq\llbracket 1,n-1\rrbracket$. First observe that $H$ cannot contain $n+1$, otherwise $a_{H}\geq a_{n+1}=a_{\llbracket 1,n\rrbracket}+a_{n}>a_{\llbracket 1,n\rrbracket}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{\llbracket 1,n\rrbracket}+a_{I}$, so we are left to check only the first two of the above conditions. If $H\subseteq\llbracket 1,n-1\rrbracket$, Lemma 3.8 implies $\llbracket 1,n\rrbracket\cap I=\emptyset$ and this can happen only if $I=\emptyset$. If $H=\llbracket 1,n\rrbracket$, then $a_{I}=0$ and $I=\emptyset$.
(v) Firstly, if $I\cap J=\emptyset$, then $a_{I}+a_{J}=a_{I\cup J}\in C_{n}$, while if $I\cup J=\llbracket 1,n\rrbracket$ and $n\in I\cap J$, then $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}=a_{\llbracket 1,n\rrbracket}+a_{n}+a_{((I\cap J)\setminus\{n\})}=a_{((I\cap J)\setminus\{n\})\cup\{n+1\}}\in C_{n}$.
On the other hand, assume $a_{I}+a_{J}\in C_{n}$, i.e., there exists $H\subseteq\llbracket 1,n+1\rrbracket$ such that
$a_{H}=a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}=a_{(I\cup J)\setminus(I\cap J)}+2a_{I\cap J}.$ (6)
If $H=\emptyset$, then $I,J=\emptyset$ and $I\cup J=\emptyset$. Assume then $H\neq\emptyset$, implying $I\cup J\neq\emptyset$. Set $h:=\max H$ and $h^{\prime}:=\max I\cup J$. We first observe that it must be $h\geq h^{\prime}$. In fact, if $h<h^{\prime}$ and $h^{\prime}\leq n$, then $a_{I}+a_{J}\geq a_{h^{\prime}}>2a_{\llbracket 1,h\rrbracket}\geq a_{H}$, while if $h^{\prime}=n+1$ (and so $h\leq n$), then $a_{I}+a_{J}\geq a_{n+1}=a_{\llbracket 1,n\rrbracket}+a_{n}>a_{\llbracket 1,n\rrbracket}\geq a_{H}$. Both cases contradict the assumption.
If $h=h^{\prime}$, we distinguish two sub-cases. If $h=h^{\prime}\leq n$, then $I,J,H\subseteq\llbracket 1,n\rrbracket$ and we get from Lemma 3.8 (with $k=n$) and (6) that $I\cap J=\emptyset$. If $h=h^{\prime}=n+1$, then $n+1\notin I\cap J$ (otherwise $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}\geq 2a_{n+1}>a_{H}$). Thus, (6) becomes $a_{H\setminus\{n+1\}}=a_{(I\cup J)\setminus\{n+1\}}+a_{I\cap J}$ and we get from Lemma 3.8 with $k=n$, that $(I\cup J)\setminus\{n+1\}$ and $I\cap J$ have no elements in common, i.e., $I\cap J=\emptyset$.
It remains to consider the case $h>h^{\prime}$. This can occur only if $h=n+1$ and $h^{\prime}=n$. In fact, if $h=n+1$ and $h^{\prime}\leq n-1$, then $a_{H}\geq a_{n+1}>a_{n}>2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$, while if $h\leq n$, then $a_{H}\geq a_{h}>2a_{\llbracket 1,h-1\rrbracket}\geq a_{I}+a_{J}$. Moreover, it must be $n\in I\cap J$, otherwise we get the following contradiction: $a_{H}\geq a_{n+1}>2a_{n}>a_{n}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$. In addition, we have $n\notin H$, otherwise $a_{H}\geq a_{n}+a_{n+1}=2a_{n}+a_{\llbracket 1,n\rrbracket}>2a_{n}+3a_{\llbracket 1,n-1\rrbracket}\geq 2a_{\llbracket 1,n\rrbracket}\geq a_{I}+a_{J}$, which is again an absurd. Finally, summing up all the observations, we have $\max H=n+1$, $\max I\cup J=n$, $H\setminus\{n+1\}\subseteq\llbracket 1,n-1\rrbracket$, $I\cup J\subseteq\llbracket 1,n\rrbracket$ and $n\in I\cap J$. Then, (6) becomes $a_{H}=a_{H\setminus\{n+1\}}+a_{n+1}=a_{H\setminus\{n+1\}}+a_{\llbracket 1,n-1\rrbracket}+2a_{n}=a_{I\setminus\{n\}}+a_{J\setminus\{n\}}+2a_{n}=a_{I}+a_{J}$, and we get
$a_{H\setminus\{n+1\}}+a_{\llbracket 1,n-1\rrbracket}=a_{I\setminus\{n\}}+a_{J\setminus\{n\}}.$
It then follows from Lemma 3.9 that $(I\setminus\{n\})\cup(J\setminus\{n\})=\llbracket 1,n-1\rrbracket$, thus $I\cup J=\llbracket 1,n\rrbracket$ and we already know that $n\in I\cap J$. All possible cases have been examined. ∎
We are finally ready to prove Theorem 3.7.
Proof of Theorem 3.7.
(1) As ${\mathcal{I}}(R)$ is unit-cancellative and reduced, to prove that $I_{B_{n}}$ is an atom of ${\mathcal{I}}(R)$ it suffices to show that there cannot be $\mathfrak{a},\mathfrak{b}\in{\mathcal{I}}(R)$ such that $I_{B_{n}}\subsetneq\mathfrak{a},\mathfrak{b}\subsetneq R$, and $I_{B_{n}}=\mathfrak{a}\cdot\mathfrak{b}$. Let us assume to the contrary that such $\mathfrak{a}$ and $\mathfrak{b}$ exist. Set $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$ and $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$. Then, by Lemma 3.1, we have $d\geq 1$, $e\geq 1$, $d+e=m=\max(B_{n})$,
$(I_{B_{n}})_{K}={\rm Span}_{K}\{X^{m-b}Y^{b}\colon b\in B_{n}\}=\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e],$ (7)
Furthermore, there exist $r\geq 2$ linearly independent elements $f_{1},\dots,f_{r}\in\mathfrak{a}_{K}[d]$ such that $X^{d}=\operatorname{\mathsf{in}}(f_{1})>\dots>\operatorname{\mathsf{in}}(f_{r})$ and $\mathfrak{a}_{K}[d]={\rm Span}_{K}\{f_{1},\ldots,f_{r}\}$, and $s\geq 2$ linearly independent elements $g_{1},\dots,g_{s}\in\mathfrak{b}_{K}[e]$ such that $X^{e}=\operatorname{\mathsf{in}}(g_{1})>\dots>\operatorname{\mathsf{in}}(g_{s})$ and $\mathfrak{b}_{K}[e]={\rm Span}_{K}\{g_{1},\ldots,g_{s}\}$, whence for each $i\in\llbracket 2,r\rrbracket$ there is some $\alpha_{i}\in\llbracket 1,d\rrbracket$ such that $\operatorname{\mathsf{in}}(f_{i})=X^{d-\alpha_{i}}Y^{\alpha_{i}}$ and for each $j\in\llbracket 2,s\rrbracket$ there is some $\beta_{j}\in\llbracket 1,e\rrbracket$ such that $\operatorname{\mathsf{in}}(g_{j})=X^{e-\beta_{j}}Y^{\beta_{j}}$.
By symmetry, we can assume without loss of generality that $d\leq m/2=a_{\llbracket 1,n\rrbracket}$ (see Remark 3.5). Moreover, $d\geq 1$ implies $e<m=a_{\llbracket 1,n-1\rrbracket\cup\{n+1\}}$.
Set $\mathcal{P}:=\mathcal{P}(\llbracket 1,n-1\rrbracket)$, $\mathcal{P}^{+}:=\mathcal{P}\setminus\{\emptyset\}$, $\mathcal{Q}:=\mathcal{P}\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}\}$ and $\mathcal{Q}^{+}:=\mathcal{Q}\setminus\{\emptyset\}$. For every fixed $i\in\llbracket 2,r\rrbracket$, $f_{i}g_{1}$ belongs to $(I_{B_{n}})_{K}[m]$, whence also $\operatorname{\mathsf{in}}(f_{i}g_{1})=X^{m-\alpha_{i}}Y^{\alpha_{i}}=\in(I_{B_{n}})_{K}[m]$. Since $1\leq\alpha_{i}\leq d\leq a_{\llbracket 1,n\rrbracket}$, by (7) it follows that $\alpha_{i}\in B_{n}\cap\llbracket 1,a_{\llbracket 1,n\rrbracket}\rrbracket=A_{n}\cup\{a_{\llbracket 1,n\rrbracket}\}\setminus{\{0\}}$, so we can write $\alpha_{i}=a_{I}$ for some $I\in\mathcal{P}^{+}\cup\{\llbracket 1,n\rrbracket\}$. Likewise, by looking at $\operatorname{\mathsf{in}}(f_{1}g_{j})$ with $j\in\llbracket 2,s\rrbracket$, we conclude that $\beta_{j}\in B_{n}\setminus{\{0,m\}}$, and so we can write $\beta_{j}=a_{J}$ for some $J\in\mathcal{Q}^{+}\cup\{\llbracket 1,n\rrbracket\}\setminus{\{\llbracket 1,n-1\rrbracket\cup\{n+1\}\}}$. Thus, $\operatorname{\mathsf{in}}(f_{i}g_{j})=X^{m-(a_{I}+a_{J})}Y^{a_{I}+a_{J}}$ and by (7) we conclude $a_{I}+a_{J}\in B_{n}\setminus{\{0\}}$.
Assume that for some $i\in\llbracket 2,r\rrbracket$, $\operatorname{\mathsf{in}}(f_{i})=Y^{a_{\llbracket 1,n\rrbracket}}$, i.e. $I=\llbracket 1,n\rrbracket$. This can only happen if $d=e=m/2=a_{\llbracket 1,n\rrbracket}$. For some $j\in\llbracket 2,s\rrbracket$, consider $\operatorname{\mathsf{in}}(g_{j})=X^{e-a_{J}}Y^{a_{J}}$, with $a_{J}\geq e$. If $a_{J}<e=a_{\llbracket 1,n\rrbracket}$, then $J\in\mathcal{P}^{+}$ and, by Lemma 3.10(iv), $a_{J}+a_{\llbracket 1,n\rrbracket}\notin B_{n}$, implying $f_{i}g_{j}\notin(I_{B_{n}})_{K}[m]$. Hence, the only possibility is $a_{J}=e$, whence we must have $\dim\mathfrak{b}_{K}[e]=2$. An analogous argument shows that also $\dim\mathfrak{a}_{K}[d]=2$, and so $\dim(\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e])\leq 4$, contradicting (7). In fact $\dim(I_{B_{n}})_{K}[m]=|B_{n}|=2^{n}+1\geq 5$.
On the other hand, assume that for some $j\in\llbracket 2,s\rrbracket$, $\operatorname{\mathsf{in}}(g_{j})=X^{e-a_{\llbracket 1,n\rrbracket}}Y^{a_{\llbracket 1,n\rrbracket}}$, i.e. $J=\llbracket 1,n\rrbracket$. For any $I\in\mathcal{P}^{+}\cup\{\llbracket 1,n\rrbracket\}$ we have $a_{I}+a_{\llbracket 1,n\rrbracket}\notin B_{n}$ by Lemma 3.10(iv), and thus $\operatorname{\mathsf{in}}(f_{i}g_{j_{0}})\notin(I_{B_{n}})_{K}[m]$ for every $i\in\llbracket 2,r\rrbracket$. It follows that $\dim\mathfrak{a}_{K}[d]=r=1$, but this is impossible.
Therefore, we can restrict our attention to those $\alpha_{i}=a_{I}$ and $\beta_{j}=a_{J}$ with $I\in\mathcal{P}^{+}$ and $J\in\mathcal{Q}^{+}\setminus{\{\llbracket 1,n-1\rrbracket\cup\{n+1\}\}}$. Furthermore, of all such possible choices of $I$ and $J$, by Lemma 3.10(iv), we need only consider those for which $I\cap J=\emptyset$.
Since $\operatorname{\mathsf{in}}(f_{i_{1}})\neq\operatorname{\mathsf{in}}(f_{i_{2}})$ whenever $i_{1}\neq i_{2}\in\llbracket 1,d\rrbracket$, and $|\mathcal{P}|=2^{n-1}$, we obtain $r=\dim\mathfrak{a}_{K}[d]\leq 2^{n-1}$. Assume that $r\in\llbracket 2^{k}+1,2^{k+1}\rrbracket$, where $k\in\llbracket 0,n-2\rrbracket$. To each $f_{i}$, where $i\in\llbracket 1,r\rrbracket$, corresponds a set $I_{i}\in\mathcal{P}$, and $I_{i_{1}}\neq I_{i_{2}}$ for each $i_{1}\neq i_{2}$. Since $r\geq 2^{k}+1$, the union of all the $I_{i}$’s has cardinality at least equal to $k+1$, since a set of cardinality at most $k$ has no more than $2^{k}$ subsets. Thus, the number of sets $J\in\mathcal{P}$ such that $J\cap I_{i}=\emptyset$ for every $i$, is at most $2^{n-k-2}$ ($J$ is actually a subset of $\llbracket 1,n-1\rrbracket\setminus\bigcup I_{i}$). Similarly, also the number of admissible sets $J\in\mathcal{Q}\setminus\mathcal{P}$ is at most $2^{n-k-2}$. In total, the number of possibilities for $J\in\mathcal{Q}$ is at most $2^{n-k-1}$. Hence $\dim\mathfrak{b}_{K}[e]\leq 2^{n-k-1}$ and consequently $\dim(\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e])\leq 2^{n}$, a contradiction to (7).
(2) $\{2,n+1\}\subseteq\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})$ follows by Remark 3.6, the previous point (1) and the fact that all the ideals of the form $\mathfrak{b}_{i}$ are atoms of ${\mathcal{I}}(R)$ (by [21, Prop. 5.10.1]).
For the proof of $\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})\subseteq\llbracket 2,n+1\rrbracket$, we assume to the contrary that the following identity holds
$I_{C_{n}}=\prod_{i=1}^{n+2}I_{i},$
where $I_{i}\in\mathcal{I}(R)\setminus\{R\}$ with $\operatorname{\mathsf{m-deg}}(I_{i})=d_{i}$ for every $i\in\llbracket 1,n+2\rrbracket$. We have $\sum_{i=1}^{n+2}d_{i}=a_{\llbracket 1,n+1\rrbracket}$ by (1). Thus,
$(I_{C_{n}})_{K}[\mu]={\rm Span}_{K}\{X^{\mu-c}Y^{c}\colon c\in C_{n}\}=\prod_{i=1}^{n+2}(I_{i})_{K}[d_{i}],$
where $\mu=a_{\llbracket 1,n+1\rrbracket}$. As in the previous point, we may assume that
$d_{i}\geq 1\text{ and }\dim(I_{i})_{K}[d_{i}]\geq 2\text{ for all }i\in\llbracket 1,n+2\rrbracket.$
Moreover, let
$(I_{i})_{K}[d_{i}]={\rm Span}\{f_{1}^{(i)},\ldots,f_{r_{i}}^{(i)}\},$
where $f_{1}^{(i)},\ldots,f_{r_{i}}^{(i)}$ are linearly independent, $\operatorname{\mathsf{in}}(f_{1}^{(i)})>\operatorname{\mathsf{in}}(f_{2}^{(i)})>\ldots>\operatorname{\mathsf{in}}(f_{r_{i}}^{(i)})$ and $\operatorname{\mathsf{in}}(f_{1}^{(i)})=X^{d_{i}}$, for all $i\in\llbracket 1,n+2\rrbracket$. For any $j\in\llbracket 2,r_{i}\rrbracket$, let $\operatorname{\mathsf{in}}(f_{j}^{(i)})=X^{d_{i}-a_{ij}}Y^{a_{ij}}$ for some $1\leq a_{ij}\leq d_{i}$. Then the following equation
$\operatorname{\mathsf{in}}\left(f_{j}^{(i)}\cdot\prod_{k=1,k\neq i}^{n+2}f_{1}^{k}\right)=X^{\mu-a_{ij}}Y^{a_{ij}}$
shows that $a_{ij}\in C_{n}$ and hence $a_{ij}=a_{I_{j}^{(i)}}$ with $\emptyset\neq I_{j}^{(i)}\subseteq\llbracket 1,n+1\rrbracket$ for every $i\in\llbracket 1,n+2\rrbracket$ and $j\in\llbracket 2,r_{i}\rrbracket$. Likewise, $\sum_{\ell=1}^{k}a_{I_{j_{\ell}}^{(i_{\ell})}}\in C_{n}$ for every $k\leq n+2$ and $i_{s}\neq i_{t}$ for every $s\neq t$. We claim that $n\in I_{2}^{(i)}$ for all $i\in\llbracket 1,n+2\rrbracket$ and that $I_{2}^{(i)}\subseteq\llbracket 1,n\rrbracket$ for all $i\in\llbracket 1,n+2\rrbracket$. If $n\not\in I_{2}^{(i_{0})}$ for some $i_{0}$, since $a_{I_{2}^{(i_{0})}}+a_{I_{2}^{(i)}}\in C_{n}$ for every $i\neq i_{0}$, by Lemma 3.10(v) we get
$I_{2}^{(i_{0})}\cap I_{2}^{(i)}=\emptyset\text{ for all }i\neq i_{0}.$
But then $|\cup_{i=1}^{n+2}I_{2}^{(i)}|\geq n+2$ (recall that $I_{2}^{(i)}\neq\emptyset$ for all $i$), contradicting the fact that $\cup_{i=1}^{n+2}I_{2}^{(i)}\subseteq\llbracket 1,n+1\rrbracket$. If we assume that $n+1\in I_{2}^{(i_{0})}$, the same argument leads to another contradiction (note that $I_{2}^{(i_{0})}\cup I_{2}^{(i)}\neq\llbracket 1,n\rrbracket$ for all $i\neq i_{0}$).
Therefore, using Lemma 3.10(v) again, we conclude
$I_{2}^{(i)}\cup I_{2}^{(j)}=\llbracket 1,n\rrbracket\text{ for all }i\neq j.$
Recall that $n\geq 2$ and $\sum_{i=1}^{n+2}a_{I_{2}^{(i)}}\in C_{n}$. Then
$\sum_{i=1}^{n+2}a_{I_{2}^{(i)}}\geq\sum_{i=1}^{4}a_{I_{2}^{(i)}}=a_{I_{2}^{(1)}\cup I_{2}^{(2)}}+a_{I_{2}^{(1)}\cap I_{2}^{(2)}}+a_{I_{2}^{(3)}\cup I_{2}^{(4)}}+a_{I_{2}^{(3)}\cap I_{2}^{(4)}},$
which is equal to
$a_{\llbracket 1,n\rrbracket}+a_{I_{2}^{(1)}\cap I_{2}^{(2)}}+a_{\llbracket 1,n\rrbracket}+a_{I_{2}^{(3)}\cap I_{2}^{(4)}}\geq 2a_{\llbracket 1,n\rrbracket}+2a_{n}=a_{\llbracket 1,n+1\rrbracket}+a_{n}>\mu,$
contradicting $\sum_{i=1}^{n+2}a_{I_{2}^{(i)}}\in C_{n}$. ∎
In [16, Prop. 4.10], the authors show that $B_{n}$ is an atom in $\mathcal{P}_{\text{fin},0}(\mathbb{N})$ and that $\mathsf{L}_{\mathcal{P}_{\text{fin},0}(\mathbb{N})}(I_{C_{n}})=\{2,n+1\}$. In light of theorem 3.7, it makes sense to ask whether, in analogy with the results in [16], also $\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})=\{2,n+1\}$; we note that this would immediately imply $\Delta({\mathcal{I}}(R))=\mathbb{N}^{+}$. The question remains open but in the next section we will prove that in the case $D=K$ the answer is negative, if we restrict the problem from $\mathcal{I}(R)$ to its submonoid ${\mathcal{M}{\rm on}}(R)$ of nonzero monomial ideals in $R$.
4. Arithmetical properties of ${\mathcal{M}{\rm on}}(R)$
In this section we focus on the arithmetical properties of the monoid ${\mathcal{M}{\rm on}}(R)$ of nonzero monomial ideals in $R=K[X_{1},\dots,X_{N}]$, with $K$ an arbitrary field and $N\geq 2$. We start by recalling a basic fact on monomial ideals which we will repeatedly use in the following discussion. We refer to [13, Cor. 1.8] for a proof.
Lemma 4.1.
Let $I\in{\mathcal{M}{\rm on}}(R)$, and $\mathcal{M}$ a set of monomials in $I$. Then $\mathcal{M}$ generates $I$ if and only if for each monomial $v\in I$ there exists $u\in\mathcal{M}$ such that $u\mid v$.
Remark 4.2.
If $I\in{\mathcal{M}{\rm on}}(R)$ and $\operatorname{\mathsf{m-deg}}{(I)}=0$, then $I=R$. Indeed, $I$ contains an element of min-degree zero, i.e., a polynomial $f\in R$ with $f(0)\neq 0$; since $I$ is generated by monomials and $f\in I$, necessarily $f(0)\in I$, whence $1\in I$.
Example 4.3.
As a first application of the lemma, we show that the ideal
$\mathfrak{c}_{4}=I_{\{0,1,2,4\}}=\langle X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}\rangle$
is an atom of ${\mathcal{M}{\rm on}}(R)$, while back in Section 2 we recalled that it is not an atom of $\mathcal{I}(R)$ (at least when the characteristic of $K$ is not $2$), despite $\{0,1,2,4\}$ being an atom of $\mathcal{P}_{\rm fin}(\mathbb{N})$. Suppose that $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$, with $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$ and $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$. By equation (1) and Remark 4.2, $d+e=4$, $d\geq 1$ and $e\geq 1$. By symmetry, we can also assume $d\leq e$, so the only possibilities for $(d,e)$ are $(1,3)$ and $(2,2)$.
Case $(d,e)=(1,3)$: first, we claim that $Y^{3}\in\mathfrak{b}$. Indeed, if this is not the case, for every monomial $X^{i}Y^{j}\in\mathfrak{b}$ we have either $i\geq 1$ or $j\geq 4$, as $e=3$; the product of such elements with monomials of $\mathfrak{a}$, all of which have positive min-degree, yields elements of the form $X^{k}Y^{l}\in\mathfrak{a}\cdot\mathfrak{b}$, with either $k\geq 1$ or $l\geq 5$. In particular, $Y^{4}\not\in\mathfrak{a}\cdot\mathfrak{b}$, a contradiction. Now suppose $X\notin\mathfrak{a}$; since $d=1$, for every monomial $X^{i}Y^{j}\in\mathfrak{a}$ we have either $i\geq 2$ or $j\geq 1$. The product of such elements with monomials of $\mathfrak{b}$ yields elements of the form $X^{k}Y^{l}$, with either $k+l\geq 5$ or $l\geq 1$. In particular, $X^{4}\not\in\mathfrak{a}\cdot\mathfrak{b}$, again a contradiction; therefore, $X\in\mathfrak{a}$. We then obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, absurd by Lemma 4.1, as none of the monomials $X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}$ divide $XY^{3}$, despite forming a generating set of $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$.
Case $(d,e)=(2,2)$: by arguments analogous to those used in the previous case, we obtain $X^{2},Y^{2}\in\mathfrak{a}\cap\mathfrak{b}$. Since $\langle X^{2},Y^{2}\rangle^{2}=\langle X^{4},X^{2}Y^{2},Y^{4}\rangle\subsetneq\mathfrak{c}_{4}$, we may assume without loss of generality that $\langle X^{2},Y^{2}\rangle\subsetneq\mathfrak{a}$. By Lemma 4.1, there exist nonnegative integers $i$ and $j$ such that $X^{i}Y^{j}\in\mathfrak{a}$, and neither $X^{2}$ nor $Y^{2}$ divides $X^{i}Y^{j}$; since $\operatorname{\mathsf{m-deg}}(X^{i}Y^{j})\geq d=2$, the only possibility is $(i,j)=(1,1)$, i.e., $XY\in\mathfrak{a}$. Multiplying with $Y^{2}\in\mathfrak{b}$, we obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, a contradiction (as in the previous case).
As we have found a contradiction in all the possible cases, we conclude that $\mathfrak{c}_{4}$ does not admit a factorization as the product of two nonunit monomial ideals.
Remark 4.4.
An important consequence of Lemma 4.1 is that every monomial ideal II admits a unique generating set of monomials that is minimal with respect to divisibility [13, Prop. 1.11]; we denote this set by $G(I)$. As a consequence of Dickson’s Lemma [13, Theorem 1.9], $G(I)$ is finite for every $I\in{\mathcal{M}{\rm on}}(R)$.
Next, we see to what extent can [21, Theorem 5.1] be adapted to the monoid ${\mathcal{M}{\rm on}}(R)$. Let us define a further class of monomial ideals, namely $\mathfrak{a}_{k}={\langle X,Y\rangle}^{k}$, for any $k\in\mathbb{N}^{+}$.
Theorem 4.5.
The monoid of monomial ideals ${\mathcal{M}{\rm on}}(R)$ of the ring $R=K[X_{1}\dots,X_{N}]$, where $K$ is a field and $N\geq 2$, is a BF-monoid. Moreover,
  1. (1)
    ${\mathcal{M}{\rm on}}(R)$ is neither locally finitely generated nor transfer Krull;
  2. (2)
    $\mathcal{U}_{k}({\mathcal{M}{\rm on}}(R))=\mathbb{N}_{\geq 2}$ for all $k\geq 2$;
  3. (3)
    $\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{k})=\llbracket 2,k\rrbracket$ for all $k\geq 2$;
  4. (4)
    ${\mathcal{M}{\rm on}}(R)$ is fully elastic.
Proof.
Let $I\in{\mathcal{M}{\rm on}}(R)$ with $m=\operatorname{\mathsf{m-deg}}{(I)}$. Firstly, since every nonunit (in particular, every atom) of ${\mathcal{M}{\rm on}}(R)$ has positive min-degree by Remark 4.2, equation (1) implies that the length of any factorization (into atoms of ${\mathcal{M}{\rm on}}(R)$) of $I$ is at most $m$. We will now show by induction on $m$ that $I$ always admits at least one factorization, thus proving that ${\mathcal{M}{\rm on}}(R)$ is a BF-monoid.
If $m=0$, then $I=R$, by Remark 4.2. When $m\in\mathbb{N}^{+}$, either $I$ is itself an atom, or there exist $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$ such that $I=\mathfrak{a}\cdot\mathfrak{b}$. By equation (1) and Remark 4.2, we have $m=\operatorname{\mathsf{m-deg}}{(\mathfrak{a})}+\operatorname{\mathsf{m-deg}}{(\mathfrak{b})}$, $\operatorname{\mathsf{m-deg}}{(\mathfrak{a})}\geq 1$ and $\operatorname{\mathsf{m-deg}}{(\mathfrak{b})}\geq 1$. Thus, $\operatorname{\mathsf{m-deg}}{(\mathfrak{a})}<m$ and $\operatorname{\mathsf{m-deg}}{(\mathfrak{b})}<m$, whence by the inductive hypothesis both $a$ and $b$ can be factorized into atoms.
(1) The fact that ${\mathcal{M}{\rm on}}(R)$ is not locally finitely generated follows from the proof of the corresponding statement for ${\mathcal{I}}(R)$ in [21, Lemma 5.3.2], since all the ideals involved are monomial, and a monomial ideal which is an atom in ${\mathcal{I}}(R)$ is clearly also an atom in ${\mathcal{M}{\rm on}}(R)$. For the second part of the statement, assume to the contrary that there are a Krull monoid $H$ and a transfer homomorphism $\theta\colon{\mathcal{M}{\rm on}}(R)\to H$. Consider the monomial ideals
$\mathfrak{a}_{1}=\langle X,Y\rangle,\ \mathfrak{a}_{2}={\langle X,Y\rangle}^{2}={\mathfrak{a}_{1}}^{2},\ \mathfrak{b}_{2}=\langle X^{2},Y^{2}\rangle.$
A straightforward computation shows that $\mathfrak{a}_{1}\mathfrak{a}_{2}=\mathfrak{a}_{1}\mathfrak{b}_{2}$, whence $\theta(\mathfrak{a}_{1})\theta(\mathfrak{a}_{2})=\theta(\mathfrak{a}_{1})\theta(\mathfrak{b}_{2})$ in $H$. Since Krull monoids are cancellative, we infer $\theta(\mathfrak{a}_{2})=\theta(\mathfrak{b}_{2})$. By Lemma 2.1, for any $I\in{\mathcal{M}{\rm on}}(R)$ we have that $I\in\mathcal{A}({\mathcal{M}{\rm on}}(R))$ if and only if $\theta(I)\in\mathcal{A}(H)$. Since $\mathfrak{b}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, it follows that $\theta(\mathfrak{b}_{2})$ is an atom of $H$. Thus, $\theta(\mathfrak{a}_{2})$ is also an atom of $H$, whence $\mathfrak{a}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, a contradiction.
(2) One can proceed exactly as in [21, Lemma 5.3.3] to show that $\mathcal{U}_{2}({\mathcal{M}{\rm on}}(R))=\mathbb{N}_{\geq 2}$, as all the ideals occurring in that proof are monomial. Since ${\mathcal{M}{\rm on}}(R)$ is a BF-monoid, the conclusion then follows by [21, Lemma 3.1.3].
(3) Clearly, $\operatorname{\mathsf{m-deg}}{(\mathfrak{a}_{k})}=k$ and, for $k\geq 2$, the ideal $\mathfrak{a}_{k}$ is not an atom of ${\mathcal{M}{\rm on}}(R)$. Since the length of a factorization of a monomial ideal is bounded by its min-degree and $1\in L_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{k})$ if and only if $\mathfrak{a}_{k}$ is an atom (i.e., if and only if $k=1$), we infer that $L_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{k})\subseteq\llbracket 2,k\rrbracket$, for any $k\geq 2$. For the converse inclusion we refer to the proof of [21, Theorem 5.1.3], noting that all the ideals therein are monomial, with one exception: when proving that $2\in L_{{\mathcal{I}}(R)}(\mathfrak{a}_{5})$, the authors use a factorization into genuine polynomial ideals to get around the fact that $\mathfrak{c}_{4}$ is generally not an atom of ${\mathcal{I}}(R)$. In our case, by considering the identity
$\mathfrak{a}_{5}={\langle X,Y\rangle}^{5}=\langle X,Y\rangle\langle X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}\rangle=\mathfrak{a}_{1}\mathfrak{c}_{4},$
involving only monomial ideals, and recalling Remark 4.3, we can conclude that $2\in L_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{5})$.
(4) Set $\mathfrak{p}=\langle X\rangle$, denote by $H_{1}$ the free monoid with basis $\{\mathfrak{p}\}$ and let $H_{2}=\{\mathfrak{a}\in{\mathcal{M}{\rm on}}(R)\colon\mathfrak{p}\nmid\mathfrak{a}\}$. The ideal $\mathfrak{p}$ is principal, thus invertible in the monoid of nonzero fractional ideals of $R$. In particular, it is a cancellative element of this monoid, hence also a cancellative element of ${\mathcal{M}{\rm on}}(R)$. Moreover, $\mathfrak{p}$ is clearly prime, whence we obtain ${\mathcal{M}{\rm on}}(R)=H_{1}\times H_{2}$. Now observe that $\mathfrak{a}_{k}\in H_{2}$ for all $k\in\mathbb{N}^{+}$, as $Y^{k}\in\mathfrak{a}_{k}$. By item (3), we can then apply [21, Prop. 3.2] to conclude that ${\mathcal{M}{\rm on}}(R)$ is fully elastic. ∎
4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
We continue our investigation of ${\mathcal{M}{\rm on}}(R)$ by constructing new classes of atoms in this monoid. Fix a positive integer $n\geq 3$ and $a_{1},\dots,a_{n},a_{n+1}\in\mathbb{N}^{+}$ verifying conditions (C1), $a_{n+1}=a_{1}+\dots+a_{n-1}+2a_{n}$, and (C2), $a_{i+1}>2(a_{1}+\dots+a_{i})$ for all $i\in\llbracket 1,n-1\rrbracket$. Keeping the notation of Section 3.2, for every nonempty $I\subseteq\llbracket 1,n+1\rrbracket$ we set $a_{I}=\sum_{i\in I}a_{i}$ while we set $a_{\emptyset}=0$. For $r\in\llbracket 3,n\rrbracket$ let us define:
$\tilde{\mathfrak{b}}_{r}:=\langle\mathfrak{b}_{a_{1}}\cdot\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}},X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\rangle.$ (8)
Note that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ is a proper generator of $\tilde{\mathfrak{b}}_{r}$, i.e., $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\notin\mathfrak{b}_{a_{1}}\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}}$. To show that this is true, we first observe that
$\mathfrak{b}_{a_{1}}\cdot\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}}=\langle X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\colon J\subseteq\{1\}\cup\llbracket 3,r\rrbracket\rangle,$
and, by Lemma 4.1, if $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\mathfrak{b}_{a_{1}}\cdot\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}}$, one of the generators of $\mathfrak{b}_{a_{1}}\cdot\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}}$ must divide it. Observe that for $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$
$a_{J}>a_{1}\Longrightarrow a_{J}\geq a_{3}.$ (9)
Thus, if there exists $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ such that $X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}$ divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, then it must be $a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}\leq a_{\llbracket 3,r\rrbracket}-a_{2}$, which implies $a_{J}\geq a_{1}+a_{2}$ and, by (9), $a_{J}\geq a_{3}$. Moreover, we must also have $a_{J}\leq a_{3}-a_{2}$, but this contradicts the previous inequality.
Proposition 4.6.
Let $n\geq 3$ be a positive integer. The following hold for every $r\in\llbracket 3,n\rrbracket$:
(5) $\mathfrak{b}_{a_{2}}\cdot\tilde{\mathfrak{b}}_{r}=\prod_{i=1}^{r}\mathfrak{b}_{a_{i}}$.
(6) $I_{C_{n}}=\mathfrak{b}_{a_{2}}\cdot\tilde{\mathfrak{b}}_{r}\cdot\prod_{i=r+1}^{n+1}\mathfrak{b}_{a_{i}}$.
Proof.
(5) By definition $\mathfrak{b}_{a_{i}}=\langle X^{a_{i}},Y^{a_{i}}\rangle$ and $\tilde{\mathfrak{b}}_{r}=\langle\mathfrak{b}_{a_{1}}\cdot\mathfrak{b}_{a_{3}}\cdots\mathfrak{b}_{a_{r}},X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\rangle$, whence
$\mathfrak{b}_{a_{2}}\cdot\tilde{\mathfrak{b}}_{r}=\langle\prod_{i=1}^{r}\mathfrak{b}_{a_{i}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{3}-a_{2}},X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}}\rangle.$
It is easy to see that $X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}+a_{2}},X^{a_{\llbracket 1,r\rrbracket}-a_{3}}Y^{a_{3}}\in\prod_{i=1}^{r}\mathfrak{b}_{a_{i}}$. But then
$X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{3}-a_{2}}=X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}+a_{2}}\cdot Y^{a_{3}-2a_{2}-a_{1}}\in\prod_{i=1}^{r}\mathfrak{b}_{a_{i}},$
since $a_{3}-2a_{2}-a_{1}>0$ by condition (C2), and
$X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}}=X^{a_{\llbracket 1,r\rrbracket}-a_{3}}Y^{a_{3}}\cdot X^{a_{3}-2a_{2}-a_{1}}\in\prod_{i=1}^{r}\mathfrak{b}_{a_{i}}.$
Therefore, $\mathfrak{b}_{a_{2}}\cdot\tilde{\mathfrak{b}}_{r}=\prod_{i=1}^{r}\mathfrak{b}_{a_{i}}$ as claimed.
(6) It follows from item (5) and Remark 3.6. ∎
The relevance of the above proposition is made clear by the following result, whose proof is provided in Section 4.2.
Theorem 4.7.
Let $n\geq 3$ be a fixed integer. For every $r\in\llbracket 3,n\rrbracket$ ($n\geq 3$), $\tilde{\mathfrak{b}}_{r}$ is an atom of ${\mathcal{M}{\rm on}}(R)$.
The construction of the class of atoms $\tilde{\mathfrak{b}}_{r}$ of ${\mathcal{M}{\rm on}}(R)$ gives us the opportunity to answer negatively, for this monoid, the question posed ad the end of Section 3.
Corollary 4.8.
For every fixed integer $n\geq 2$, $\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})=\llbracket 2,n+1\rrbracket$.
Proof.
We already know by Theorem 3.7(2) that $\{2,n+1\}\subseteq\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})\subseteq\llbracket 2,n+1\rrbracket$ for every $n\geq 2$. If $n=2$ the conclusion is trivial; if $n\geq 3$ it follows from Proposition 4.6(6) and Theorem 4.7. ∎
We conclude with a further theorem on the atomicity of $2$-generated ideals $\langle X^{n},Y^{m}\rangle$. The proof is given in Section 4.2.
Theorem 4.9.
For every $m,n\in\mathbb{N}^{+}$, the ideal $I=\langle X^{m},Y^{n}\rangle$ is an atom in ${\mathcal{M}{\rm on}}(R)$.
4.2. Proofs of Theorems 4.7 and 4.9
We start our discussion by proving some technical lemmas.
Lemma 4.10.
Fix a positive integer $n\geq 4$. Let $I,J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ with $r\in\llbracket 4,n\rrbracket$. If $a_{J}-a_{I}>a_{1}$ and $a_{J}-a_{I}\neq a_{3}-a_{1}$, then $a_{J}-a_{I}\geq a_{3}$.
Proof.
We first observe that $a_{J}-a_{I}=a_{J\setminus I}-a_{I\setminus J}$. If $I\setminus J=\emptyset$, then $a_{J}-a_{I}=a_{J\setminus I}$ and $a_{J\setminus I}>a_{1}$ trivially implies $a_{J\setminus I}\geq a_{3}$, as $J\setminus I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$. Assume then $I\setminus J\neq\emptyset$. Clearly our assumptions imply that also $J\setminus I\neq\emptyset$ (otherwise $a_{J}-a_{I}<0$), so there exist $s,t>0$ such that $s=\max(J\setminus I)$ and $t=\max(I\setminus J)$. Note that, since $(J\setminus I)\cap(I\setminus J)=\emptyset$, $s\neq t$. Moreover, it must be $s>t$. In fact, if $s<t$ (implying that $t\geq 3$), then
$a_{I\setminus J}\geq a_{t}>a_{\llbracket 1,t-1\rrbracket}\geq a_{\llbracket 1,s\rrbracket}\geq a_{J\setminus I},$
but this is absurd since we are assuming $a_{J}-a_{I}=a_{J\setminus I}-a_{I\setminus J}>0$. We can also prove that $s\geq 4$. If this is not the case, in fact, the only admissible alternative is $s=3$ and $t=1$. However, this leads to $I\setminus J=\{1\}$ and $J\setminus I=\{3\}$, i.e., to $a_{J}-a_{I}=a_{3}-a_{1}$, a case that we discarded by assumption.
Therefore, we can conclude that
$a_{J}-a_{I}=a_{J\setminus I}-a_{I\setminus J}\geq a_{s}-a_{\llbracket 1,s-1\rrbracket}>a_{\llbracket 1,s-1\rrbracket}\geq a_{\llbracket 1,3\rrbracket}>a_{3}.$
In the following proofs, we will use without mention the following property of minimal generating sets of products of monomial ideals, whose proof is straightforward: given $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$, we have
$G(\mathfrak{a}\cdot\mathfrak{b})\subseteq\{fg\colon f\in G(\mathfrak{a}),g\in G(\mathfrak{b})\}.$
Lemma 4.11.
Let $I,\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus{\{R\}}$ with $\operatorname{\mathsf{m-deg}}(I)=m$, $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$, $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$, $X^{m},Y^{m}\in I$ and $I=\mathfrak{a}\cdot\mathfrak{b}$. Then:
(7) $d,e\geq 1$.
(8) $X^{d},Y^{d}\in G(\mathfrak{a})$and $X^{e},Y^{e}\in G(\mathfrak{b})$.
Proof.
(7) Apply Remark 4.2.
(8) First we prove that $X^{d},Y^{d}\in\mathfrak{a}$ and $X^{e},Y^{e}\in\mathfrak{b}$. By equation (1), $m=d+e$. Now if $X^{i}\not\in\mathfrak{a}$ for every $i\in\llbracket 1,m\rrbracket$, then $X^{m}\not\in I$. Therefore $X^{i_{0}}\in\mathfrak{a}$ for some $i_{0}\in\llbracket 1,m\rrbracket$ such that $i_{0}\geq d$ and hence $X^{m-i_{0}}\in\mathfrak{b}$, where $m-i_{0}\geq e$. If $i_{0}>d$, then $i_{0}+e>d+e=m$ and hence $e>m-i_{0}$, which contradicts the fact that $\operatorname{\mathsf{m-deg}}{(\mathfrak{b})}=e$. So $i_{0}=d$ and hence $X^{d}\in\mathfrak{a}$ and $X^{e}\in\mathfrak{b}$. Likewise, we can show that $Y^{d}\in\mathfrak{a}$ and $Y^{e}\in\mathfrak{b}$. Since $d=\operatorname{\mathsf{m-deg}}{(\mathfrak{a})}$, $X^{d},Y^{d}\in\mathfrak{a}$ and $\mathfrak{a}$ is a monomial ideal, by Lemma 4.1 we can conclude that $X^{d},Y^{d}\in G(\mathfrak{a})$; analogously, $X^{e},Y^{e}\in G(\mathfrak{b})$. ∎
Let $n\geq 3$ and $r\in\llbracket 3,n\rrbracket$. Recall that
$\tilde{\mathfrak{b}}_{r}=\langle\left\{X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\colon J\subseteq\{1\}\cup\llbracket 3,r\rrbracket\right\},X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\rangle.$ (10)
Note that $\operatorname{\mathsf{m-deg}}(\tilde{\mathfrak{b}}_{r})=a_{1}+a_{\llbracket 3,r\rrbracket}=a_{\{1\}\cup\llbracket 3,r\rrbracket}$. In fact, by definition of the $a_{i}$’s, $a_{1}<a_{3}-2a_{2}$ and so $a_{1}+a_{\llbracket 3,r\rrbracket}<a_{\llbracket 3,r\rrbracket}+a_{3}-2a_{2}$.
Lemma 4.12.
Assume $\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$ and let $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})$. Then:
(9) $\alpha,\beta\geq 1$and $a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta$;
(10) $X^{\alpha},Y^{\alpha}\in G(\mathfrak{a})$and $X^{\beta},Y^{\beta}\in G(\mathfrak{b})$;
(11) $\alpha=a_{I_{1}}$, $\beta=a_{I_{2}}$ with $I_{1},I_{2}\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and $I_{1}\cap I_{2}=\emptyset$.
Proof.
(9) Follows from Equation (1) and Lemma 4.11(7).
(10) Follows from Lemma 4.11(8).
(11) By Lemma 2.5(ii), we get $(\tilde{\mathfrak{b}}_{r})_{K}[a_{\{1\}\cup\llbracket 3,r\rrbracket}]=\mathfrak{a}_{K}[\alpha]\cdot\mathfrak{b}_{K}[\beta]$, where
$(\tilde{\mathfrak{b}}_{r})_{K}[a_{\{1\}\cup\llbracket 3,r\rrbracket}]={\rm Span}_{K}\left\{X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\colon J\subseteq\{1\}\cup\llbracket 3,r\rrbracket\right\}.$ (11)
By (10) $X^{\alpha},Y^{\alpha}\in\mathfrak{a}_{K}[\alpha]$ and $X^{\beta},Y^{\beta}\in\mathfrak{b}_{K}[\beta]$. Hence, $X^{\alpha}Y^{\beta}\in\tilde{\mathfrak{b}}_{r,K}[a_{\{1\}\cup\llbracket 3,r\rrbracket}]$, and so
$X^{\alpha}Y^{\beta}=X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{I_{2}}}Y^{a_{I_{2}}}$
for some $I_{2}\subseteq\{1\}\cup\llbracket 3,r\rrbracket$. This implies that $\alpha=a_{1}+a_{\llbracket 3,r\rrbracket}-a_{I_{2}}=a_{I_{1}}$ for some $I_{1}\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ with $I_{1}\cap I_{2}=\emptyset$ and $\beta=a_{I_{2}}$. ∎
We now have all the ingredients to finally prove Theorem 4.7.
Proof of Theorem 4.7.
Fix $n\geq 3$ and let $r\in\llbracket 3,n\rrbracket$. Assume as a contradiction that $\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$. By Lemma 4.12, $X^{\alpha},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta}\in G(\mathfrak{b})$ with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})=a_{I}$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})=a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha=a_{(\{1\}\cup\llbracket 3,r\rrbracket)\setminus I}$, where $I$ is a nonempty proper subset of $\{1\}\cup\llbracket 3,r\rrbracket$}. By symmetry, we can also assume without loss of generality that $\alpha\leq\beta$. A rough estimate ensures then that $\alpha<a_{r}$. In fact, if $\alpha\geq a_{r}$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta\geq 2\alpha\geq 2a_{r}$, which is impossible, as $a_{r}>a_{\llbracket 1,r-1\rrbracket}$. Thus, we can assume without loss of generality that, if $r\geq 4$, then $\alpha=a_{I}$ with $I\subseteq\{1\}\cup\llbracket 3,r-1\rrbracket$, while, if $r=3$, then $\alpha=a_{1}$. As a consequence, $\beta\geq a_{r}$.
It is also easy to observe that we can exclude that $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$ and $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$, i.e., that
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}}\rangle.$ (12)
In fact, the generators of $\tilde{\mathfrak{b}}_{r}$ in equation (12) are of the form $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}$, with $J=\emptyset,\{1\}\cup\llbracket 3,r\rrbracket,I,(\{1\}\cup\llbracket 3,r\rrbracket)\setminus I$. Therefore, by the observation following equation (8), none of them divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ and this leads to the contradiction $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\notin\tilde{\mathfrak{b}}_{r}$.
Now we distinguish two cases depending on whether $\alpha$ is equal or not to $a_{1}$, and for each case we consider three subcases.
CASE 1: $\alpha=a_{1}$ and $\beta=a_{\llbracket 3,r\rrbracket}$.
Subcase I: $\mathfrak{a}=\langle X^{a_{1}},Y^{a_{1}}\rangle$ and $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{l+a_{1}}Y^{m},X^{l}Y^{m+a_{1}},\dots\rangle,$ (13)
with $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$ and $1\leq l,m<a_{\llbracket 3,r\rrbracket}$.
We know from (8) that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, hence by Lemma 4.1 at least one of the generators in (13) must divide it. From remarks following (8), it can only happen that $X^{l+a_{1}}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or that $X^{l}Y^{m+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, for some $l,m$ satisfying the requirements above.
Say that there are $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$, $1\leq l,m<a_{\llbracket 3,r\rrbracket}$ and
$X^{l+a_{1}}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1},$ (14)
$m\leq a_{3}-a_{2}.$ (15)
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l}Y^{m+a_{1}}$ and this leads to the contradiction $X^{l}Y^{m+a_{1}}\notin\tilde{\mathfrak{b}}_{r}$. To prove the last assumption we first notice that, being $l<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+a_{1}}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l}Y^{m+a_{1}},$ (16)
using (14), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1}$, implying that $a_{J}\geq 2a_{1}+a_{2}$ and so that $a_{J}\geq a_{3}$ by (9). But (16) also implies $a_{J}\leq m+a_{1}$ and, by (15), we reach the contradiction $a_{J}\leq a_{3}-a_{2}+a_{1}<a_{3}$ (recall that $a_{2}>a_{1}$).
On the other hand, assume that there are $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$, $1\leq l,m<a_{\llbracket 3,r\rrbracket}$ and
$X^{l}Y^{m+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (17)
$m\leq a_{3}-a_{2}-a_{1}.$ (18)
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l+a_{1}}Y^{m}$ and this leads to the contradiction $X^{l+a_{1}}Y^{m}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last assumption we first notice that, being $m<a_{3}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l+a_{1}}Y^{m}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l+a_{1}}Y^{m},$ (19)
using (17), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l+a_{1}\leq a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{2}$, implying that $a_{J}\geq a_{2}$ and so that $a_{J}\geq a_{3}$ by (9). But (19) also implies $a_{J}\leq m$ and, by (18), we reach the contradiction $a_{J}\leq a_{3}-a_{2}-a_{1}<a_{3}$.
Subcase II: $X^{a_{1}},Y^{a_{1}},X^{i}Y^{j}\in G(\mathfrak{a})$ and $\mathfrak{b}=\langle X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}}\rangle$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},X^{i}Y^{j+a_{\llbracket 3,r\rrbracket}},\dots\rangle,$ (20)
with $i,j\in\mathbb{N}$ such that $i+j\geq a_{1}$ and $1\leq i,j<a_{1}$.
This case never occurs. In fact, for any choice of $i,j\in\mathbb{N}$ such that $i+j\geq{a_{1}}$ and $1\leq i,j<a_{1}$, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$ and this leads to the contradiction $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}\notin\tilde{\mathfrak{b}}_{r}$. To show that this last assumption is true we first notice that, being $j<a_{1}<a_{3}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},$ (21)
we get $a_{J}\leq j<a_{1}$, implying that $a_{J}=0$ (this is a trivial consequence of the definition of $J$), whence $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}>i+a_{\llbracket 3,r\rrbracket}$, a contradiction.
Subcase III: $X^{a_{1}},Y^{a_{1}},X^{i}Y^{j}\in G(\mathfrak{a})$ and $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},X^{i}Y^{j+a_{\llbracket 3,r\rrbracket}},X^{l+a_{1}}Y^{m},X^{l}Y^{m+a_{1}},\dots\rangle,$ (22)
with $i,j\in\mathbb{N}$ such that $i+j\geq a_{1}$ and $1\leq i,j<a_{1}$ and $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$ and $1\leq l,m<a_{\llbracket 3,r\rrbracket}$.
Arguing as in the previous subcase, also this case can never happen.
CASE 2: $\alpha\neq a_{1}$, i.e., $\alpha\geq a_{3}$ by Lemma 4.12(11). Under our assumption that $\alpha<a_{r}$ (and so $\beta\geq a_{r}$), this case is only possible if $r\geq 4$ (so also $n\geq 4$). As a consequence, we also have $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, where $\llbracket 4,r\rrbracket$ is well defined.
Subcase I: $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$ and $X^{\beta},Y^{\beta},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{l+\alpha}Y^{m},X^{l}Y^{m+\alpha},\dots\rangle,$ (23)
with $l,m\in\mathbb{N}$ such that $l+m\geq\beta$ and $1\leq l,m<\beta$.
We know from (8) that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, therefore by Lemma 4.1 at least one of the generators in (23) must divide it. From remarks following (8), it can only happen that $X^{l+\alpha}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or that $X^{l}Y^{m+\alpha}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, for some $l,m$ satisfying the requirements above. The latter case can be also discarded: being $\alpha\geq a_{3}$ and $m\geq 1$, $m+\alpha>a_{3}>a_{3}-a_{2}$.
We then have just one case to consider. Say that there are $l,m\in\mathbb{N}$ such that $l+m\geq\beta$, $1\leq l,m<\beta$ and
$X^{l+\alpha}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha,$ (24)
$m\leq a_{3}-a_{2}.$ (25)
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l}Y^{m+\alpha}$ and this leads to the contradiction $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last fact we first notice that, being $l<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+\alpha}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l}Y^{m+\alpha},$ (26)
using (24), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha$, implying that $a_{J}\geq a_{1}+a_{2}+\alpha$. Recalling that $\alpha=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and that $r\geq 4$, by Lemma 4.10 we have then two possibilities: $a_{J}=a_{3}-a_{1}+\alpha$ or $a_{J}\geq a_{3}+\alpha$. But (26) and (25) also imply $a_{J}\leq m+\alpha\leq a_{3}-a_{2}+\alpha$ and in both the cases above we reach a contradiction.
Subcase II: $X^{\alpha},Y^{\alpha},X^{i}Y^{j}\in G(\mathfrak{a})$ and $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+\beta}Y^{j},X^{i}Y^{j+\beta},\dots\rangle,$ (27)
with $i,j\in\mathbb{N}$ such that $i+j\geq\alpha$ and $1\leq i,j<\alpha$.
Arguing as in the previous case, there must exist $i,j\in\mathbb{N}$ such that $i+j\geq\alpha$, $1\leq i,j<\alpha$ and
$X^{i+\beta}Y^{j}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Note in fact that, since $j+\beta>a_{r}>a_{3}>a_{3}-a_{2}$, $X^{i}Y^{j+\beta}\nmid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Then,
$i\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\beta,$ (28)
$j\leq a_{3}-a_{2}.$ (29)
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{i}Y^{j+\beta}$ and this leads to the contradiction $X^{i}Y^{j+\beta}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last assumption, we first notice that, being $i<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i}Y^{j+\beta}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i}Y^{j+\beta},$ (30)
using (28), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq i\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\beta$, implying that $a_{J}\geq a_{1}+a_{2}+\beta$. Since $\beta=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and that $r\geq 4$, using Lemma 4.10, (30) and (29), we find a contradiction as in the previous subcase.
Subcase III: $X^{\alpha},Y^{\alpha},X^{i}Y^{j}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+\beta}Y^{j},X^{i}Y^{j+\beta},X^{l+\alpha}Y^{m},X^{l}Y^{m+\alpha},X^{i+l}Y^{j+m},\dots\rangle,$ (31)
with $i,j,l,m\in\mathbb{N}$ such that $i+j\geq\alpha$, $1\leq i,j<\alpha$, $l+m\geq\beta$ and $1\leq l,m<\beta$.
From (8) we know that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$ so at least one of the generators in (31) must divide it. From remarks following (8), and in view of subcases I and II, there must exist $i,j,l,m\in\mathbb{N}$ satisfying conditions above such that
$X^{i+l}Y^{j+m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Then,
$i+l\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (32)
$j+m\leq a_{3}-a_{2}.$ (33)
We claim that then, either $X^{i+\beta}Y^{j}\notin\tilde{\mathfrak{b}}_{r}$ or $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$.
To prove this assumption, let us assume that $X^{i+\beta}Y^{j}\in\tilde{\mathfrak{b}}_{r}$. In view of Lemma 4.1, one of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) must divide it. The inequality (33) implies that $j<a_{3}-a_{2}$, hence $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+\beta}Y^{j}$. Moreover, if $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ is such that $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+\beta}Y^{j}$, then $a_{J}\leq j<a_{3}-a_{2}$ and this can happen only if $a_{J}=a_{1}$ (if $a_{J}=0$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta>i+\beta$). Thus, $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\llbracket 3,r\rrbracket}\leq i+\beta$, i.e., $i\geq a_{\llbracket 3,r\rrbracket}-\beta$ and, from (32), we get $a_{\llbracket 3,r\rrbracket}-\beta+l\leq i+l\leq a_{\llbracket 3,r\rrbracket}-a_{2}$. Now, using the fact observed at the beginning of CASE 2 that $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, we obtain from the previous chain of inequalities, that $l\leq\beta-a_{2}\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}-a_{2}<a_{\llbracket 4,r\rrbracket}$ ($a_{1}-a_{2}<0$). Moreover, we derive from (33) that $m+\alpha<a_{3}-a_{2}+\alpha$.
At this point we are ready to show that $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$, thus proving our claim. First of all, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+\alpha}$ since $a_{\llbracket 3,r\rrbracket}-a_{2}>a_{\llbracket 4,r\rrbracket}>l$. Moreover, if for some $\bar{J}\subseteq\{1\}\cup\llbracket 3,r\rrbracket$
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}}Y^{a_{\bar{J}}}\mid X^{l}Y^{m+\alpha},$ (34)
then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}\leq l\leq\beta-a_{2}\leq a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha-a_{2}$, implying that $a_{\bar{J}}\geq a_{2}+\alpha$. Then, by Lemma 4.10 we have two possibilities: $a_{\bar{J}}=a_{3}-a_{1}+\alpha$ or $a_{\bar{J}}\geq a_{3}+\alpha$. But (34) also implies that $a_{\bar{J}}\leq m+\alpha<a_{3}-a_{2}+\alpha$ and we reach a contradiction in both the cases above.
All possible cases have been examined and discarded, so we can conclude that $\tilde{\mathfrak{b}}_{r}\neq\mathfrak{a}\cdot\mathfrak{b}$ for all $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$. ∎
We end this section by proving the atomicity in ${\mathcal{M}{\rm on}}(R)$ of $2$-generated ideals of the form $\langle X^{n},Y^{m}\rangle$.
Lemma 4.13.
Let $m,n\in\mathbb{N}^{+}$ such that $m>n$, and let $I\in{\mathcal{M}{\rm on}}(R)$ with $X^{m},Y^{n}\in G(I)$. Assume $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus{R}$. Then:
(12) $\operatorname{\mathsf{m-deg}}{(I)}=\alpha+\beta$, with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})\geq 1$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})\geq 1$.
(13) there exist $a,b\in\mathbb{N}^{+}$, with $a\geq\alpha$, $b\geq\beta$, and $a+b=m$ such that $X^{a},Y^{\alpha}\in\ G(\mathfrak{a})$ and $X^{b},Y^{\beta}\in\ G(\mathfrak{b})$.
Proof.
(12) This follows from equation (1) and Remark 4.2.
(13) The minimal generators of $\mathfrak{a}$ are necessarily monomials of the form $X^{a}Y^{c}$, with $a+c\geq\alpha$, while those of $\mathfrak{b}$ are monomials of the form $X^{b}Y^{d}$, with $b+d\geq\beta$. We observe that there must be a minimal generator of $\mathfrak{a}$ of the form $Y^{c}$, for some $c\geq\alpha$. Otherwise, $X$ would divide every element of $\mathfrak{a}$ and thus of $\mathfrak{a}\cdot\mathfrak{b}$, but this is impossible since $Y^{n}\in I$. Similarly, there must be a minimal generator of $\mathfrak{a}$ of the form $X^{a}$ for some positive integer $a\geq\alpha$. Arguing in the same way, $\mathfrak{b}$ also has among its minimal generators a monomial of the form $Y^{d}$ and one of the form $X^{b}$, for some $b,d\geq\beta$. Thus, $Y^{c+d}$ is a generator of $I$, while all its other minimal generators are monomials divisible by $X$. Since $Y^{n}\in I$, by Lemma 4.1, $Y^{c+d}$ must divide it, i.e, it must be $c+d\leq n$. But this can clearly happen only if $c=\alpha$ and $d=\beta$. Analogously, $X^{a+b}$ is a generator of $I$ and so it must divide $X^{m}$, whence $a+b\leq m$. But if $a+b<m$, none of the generators of $I$ divides $X^{a+b}$, so the only possibility is that $a+b=m$. ∎
Remark 4.14.
It is worth noting that for an ideal $I\in{\mathcal{M}{\rm on}}(R)$ with $X^{m},Y^{n}\in\ G(I)$ for some $m,n\in\mathbb{N}^{+}$, in order to check whether $I$ is an atom one needs only to compute a finite number of products $\mathfrak{a}\cdot\mathfrak{b}$. Indeed, by Lemma 4.13, if $I=\mathfrak{a}\cdot\mathfrak{b}$, then $X^{a},Y^{\alpha}\in\ G(\mathfrak{a})$, so every monomial $X^{c}Y^{\gamma}\in\ G(\mathfrak{a})$ satisfies $c\leq a$ and $\gamma\leq\alpha$, and analogously a monomial $X^{d}Y^{\delta}\in\ G(\mathfrak{b})$ satisfies $d\leq b$ and $\delta\leq\beta$.
Proof of Theorem 4.9.
We already know from [21, Prop. 5.10.1] that $I$ is an atom of $I(R)$, hence also of ${\mathcal{M}{\rm on}}(R)$, when $m=n$. Therefore, by symmetry, it is enough to consider the case $m>n$. If $n=1$, equation (1) ensures the atomicity of $I$ over ${\mathcal{M}{\rm on}}(R)$: if $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$, one among $\mathfrak{a}$ and $\mathfrak{b}$ has min-degree zero and must be the whole ring $R$. Let $n\geq 2$ and assume as a contradiction that there exist proper nonzero monomial ideals of $R$, say $\mathfrak{a},\mathfrak{b}$, such that $I=\mathfrak{a}\cdot\mathfrak{b}$. By Lemma 4.13, we have that $X^{a},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{b},Y^{\beta}\in G(\mathfrak{b})$, where $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})\geq 1$, $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})\geq 1$, $\alpha+\beta=n$, $a\geq\alpha\geq 1$, $b\geq\beta\geq 1$ and $a+b=m$. It follows that $X^{a}Y^{\beta}\in I$, but this is a contradiction since $m>a$ and $n>\beta$, so none of the generators of $I$ divides it. ∎
5. Prospects for future research
The results of this paper suggest several directions for future research, as they provide large families of atoms in $\mathcal{I}(R)$ and ${\mathcal{M}{\rm on}}(R)$, along with techniques to construct them. In light of the recent breakthrough by Reinhart [34] on the arithmetic of $P_{\rm fin,0}(\mathbb{N})$, we expect that evaluating finer arithmetical invariants of these monoids, such as the (monotone) catenary degree (see [22] for the definition and some examples in which it can be computed), should be feasible. Moreover, in view of Theorem 3.7(2) and Corollary 4.8, it would be interesting to attempt the construction of ideals whose sets of lengths are not intervals. Such results could, in turn, yield information about the sets of distances in the monoids under consideration. Another possible line of research would consist in detecting the strong atoms of ${\mathcal{I}}(R)$ or ${\mathcal{M}{\rm on}}(R)$; in the terminology of [14], these are atoms $a$ of a monoid $H$ for which $\mathsf{Z}_{H}(a^{n})$ is a singleton for every $n\in\mathbb{N}$. In this context, the recent paper [32] by Rath and Rissner on generating sets of powers of monomial ideals may serve as a valuable reference. Finally, we highlight the problem of computing the density of atoms in the aforementioned monoids of ideals. On this topic, it is worth recalling that in the reduced power monoid $P_{\rm fin,0}(\mathbb{N})$, almost every element is an atom [8, Theorem 6.1].